### v1

# Inferring Pluggable Types with Machine Learning

Abstract
Pluggable type systems allow programmers to extend the type system of a programming language to enforce semantic properties defined by the programmer. Pluggable type systems are difficult to deploy in legacy codebases because they require programmers to write type annotations manually. This paper investigates how to use machine learning to infer type qualifiers automatically. We propose a novel representation, NaP-AST, that encodes minimal dataflow hints for the effective inference of type qualifiers. We evaluate several model architectures for inferring type qualifiers, including Graph Transformer Network, Graph Convolutional Network and Large Language Model. We further validated these models by applying them to 12 open-source programs from a prior evaluation of the NullAway pluggable typechecker, lowering warnings in all but one unannotated project. We discovered that GTN shows the best performance, with a recall of .89 and precision of 0.6. Furthermore, we conduct a study to estimate the number of Java classes needed for good performance of the trained model. For our feasibility study, performance improved around 16k classes, and deteriorated due to overfitting around 22k classes.
Index Terms:
component, formatting, style, styling, insert
I Introduction
When programmers write code, they are often aware of semantic properties that a variable must adhere to but which are not checked by the type system of their programming language. Pluggable type systems are a method of adding custom type qualifiers to a language. A pluggable typechecker statically checks these user-defined qualifiers and warns about potential violations at compile time. For example, PUnits [1] adds qualifiers for scientific units, while Uber’s NullAway [2], Meta’s NullSafe [3], and the Checker Framework [4, 5] enforce programs do not crash due to null-pointer dereferences with a @Nullable qualifier. A key barrier to bringing legacy codebases into compliance with a typechecker is that developers must write these qualifiers into their code to see the full benefit of the tools. The state of the practice when deploying a new typechecker on a legacy codebase remains manual annotation: that is, in practice, type qualifiers are written by hand. The high cost of this manual process discourages developers from using these systems, despite their promise to improve code quality. This paper addresses the problem of bringing a legacy codebase into compliance with a typechecker by automatically inferring type qualifiers.
Recently, researchers have proposed using machine learning to solve a similar problem: inferring type annotations for untyped Python or TypeScript programs [6, 7, 8, 9, 10, 11]. This problem shares a key feature with our problem of inferring type qualifiers for a legacy codebase: in both cases, the programs were written without feedback from the typechecker, so we expect the programs not to typecheck cleanly (with any set of annotations). Experimental results suggest that these techniques outperform traditional deductive type inference techniques, so in this paper we investigate whether a similar approach can work for the problem of inferring pluggable types.
However, techniques for inferring types in Python or JavaScript programs are not immediately transferrable to our setting, because of three key differences. First, when inferring types for Python or JavaScript, the property of interest is known in advance: the type system is fixed (it enforces “tradtional” type safety: objects do not try to perform invalid operations). A key advantage of pluggable type systems is their customizability: each type system enforces its own rules. So, we require a way to customize our inference approach that does not require us to start over from scratch for each new type system. Second, for a given pluggable type system, the space of possible type qualifiers is fixed: for example, a typechecker for preventing null dereferences might only contain the qualifiers @NonNull and @Nullable. By constrast, when inferring types for e.g., Python, the space of possible types is infinite and has a long tail, because of the presence of user-defined types. That is, in the standard setting the type system is fixed and the number of types is infinite; but in the pluggable types setting there are infinitely-many possible type systems, but each has a finite, known universe of types. Third, machine learning is data hungry: approaches for popular languages like Python can assume the existence of a large corpus of human-written types in open-source software. By contrast, the customizability of pluggable types means that most type systems have few users: in fact, pluggable type systems are sometimes designed only for a single program! So, to apply machine learning techniques to the pluggable type inference setting, we need to understand how much data is actually required to make them effective. These differences inspired us to investigate three high-level research questions:
  • RQ1: What kind of ML model should we use, and how should we post-process its output?
  • RQ2: How should we encode the program in a type-system-specific way to train the model?
  • RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
The results of our investigation of these questions allows us to propose a general pipeline for training an ML model to infer type qualifiers for a given pluggable type system.
The state-of-the-art models [12, 6, 7] for inferring types in Python and JavaScript leverage the type dependence graph, a graph that encodes the type system’s semantics and which is therefore unsuitable for representing pluggable type systems because each type system has its own rules. Prior works used a variety of model architectures, including graph models and text-based models. For example, Typilus [12] learns a type space, STIR [6] uses a BiLSTM for incomplete code, and Type5 [7] uses a pre-trained seq2seq model. The differences between the pluggable types setting and the Python/JavaScript setting, however, inspired us to investigate RQ1 by evaluating a set of relatively-simple models. In particular, we eschew time-series models like BiLSTM (which are effective for sequential data, like recognizing the content of a video, but very expensive during training and inference) because we hypothesize that they may not be needed to predict pluggable type qualifiers. We considered two graph models and one text-based model. The two graph models are a baseline [13] graph convolutional network (GCN) model [14] and a more-capable graph transformer network (GTN) model [15], which extends the GCN’s capabilities by adding the ability to reason about multiple edge types (i.e., hypergraphs). The text-based model is a large language model (LLM) [16]; LLMs have performed well on coding tasks recently, but they operate in reduced context because of their token limit.
To investigate RQ2, we considered two techniques: 1) prompt the LLM about which program locations should be annotated with a type qualifier, and 2) develop a novel training methodology for building a type-system-specific graph encoding. The first technique was relatively ineffective, so we pursued the second. Our graph encoding methodology uses ablation studies to retain information that is useful for predicting the type qualifiers used by a particular type system. This process results in a novel program encoding that we call a Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), which we describe in section III. A NaP-AST encodes the parts of the program’s structure that are relevant to the type system of interest, along with name information that ties related program elements together, into a graph that is a suitable input to a graph-based model.
To evaluate these models and encodings, we conducted a feasibility study. Of the commonly used type qualifiers, those for nullability in Java are the most common. We therefore conducted our feasibility study on nullness: if there is not enough data for Java nullness type qualifiers to train an effective model, there is not enough data to train an effective model for any pluggable type system. First, we collected a dataset from Github of Java classes, each of which uses one of 31 different @Nullable annotations. After cleaning, this dataset contains 32,370 classes; more details about the dataset and its collection are presented in section IV. We then used our methodology for building a NaP-AST encoding to derive a NaP-AST encoding for nullness. Then, we trained the three models using this dataset: the two graph models (a GCN and a GTN) using the derived NaP-AST encoding, and the text-based LLM using syntax-based code chunking. We then compared their performance on 12 open-source projects that had previously been annotated for the popular NullAway typechecker [2]. The GTN outperformed the other models we evaluated: the GCN requires additional postprocessing to reach acceptable performance, while the LLM is too limited in token capacity to be effective on large projects. Our best model (the GTN) achieves 89% recall of the human-written type qualifiers and 60% precision; its inferred annotations eliminate 60% of the NullAway warnings that a human would need to triage (a proxy for human effort). Our success for nullability suggests that our approach may be effective with other type qualifiers, but unfortunately we lack enough data for any other type qualifier to test this hypothesis.
However, to assist future researchers or practitioners who would like to train a similar model for another type qualifier (and answer RQ3), we explore how much training data one needs for adequate performance. We rebuilt our best model with random samples of the dataset and re-ran the evaluation to show how much data it needs before it acheives good performance. We found that around 16k human-annotated classes were needed: a large but not impossible-to-reach number. We counted 217,922 annotations in these classes.
The primary contributions of our paper are:
  • A novel methodology to construct a per-type-system graph encoding of programs (NaP-ASTs) to train deep-learning graph models like GCNs and GTNs (section III);
  • A dataset of 32,370 Java classes annotated with @Nullable that enables our feasibility study (section IV);
  • A comparison of three model architectures (GCN, GTN, and LLM) that shows that the GTN is best for the problem of inferring pluggable types in our feasibility study (section VI-A), with open-source models and data;
  • A study of the amount of data required by our best GTN model on the task of inferring type qualifiers for adequate performance (section VI-B).
II Background: Model Architectures
This section provides background on the three considered models: graph convolutional networks (GCNs), graph transformer networks (GTNs), and large language models (LLMs). We focus on the model architectures and their salient features with respect to type qualifier inference, and also introduce the specific models that we use in our experiments.
II-A Graph Models
II-A1 GCN
Graph Convolutional Networks (GCNs) are a class of neural networks designed for graph data. Type qualifier prediction can be modeled as a graph data task because ASTs can be represented as a graph. GCNs use graph convolution, a propagation rule that updates the feature representation of each node by aggregating features from its neighbors (and the node itself). For AST prediction, GCNs can be particularly effective due to explicitly modeling local context: ASTs are inherently hierarchical and structured, where the context around a node (e.g., its parent or children nodes) significantly influences it. GCNs leverage this local structure efficiently. They are also efficient in learning spatial features, which have been helpful for ASTs [17].
A more specialized graph neural network model might outperform a GCN by adapting to heterogeneous relationships. If different types of connections between AST nodes (e.g., data flow vs. control flow) play a significant role, models within the broader GNN category designed to handle such heterogeneity might be more effective. We investigate this possibility by comparing GCN with such a model, GTN. Certain GNN architectures might also be better suited for incorporating global context or long-range dependencies within the AST, which could encode dataflow information.
II-A2 GTN
GCNs assume that the provided graphs are homogeneous, i.e., they consist of one type of nodes and edges. However, in uses cases like source code annotation distinct edge types can alter the importance of each node. Taking advantage of heterogenous graphs allows us to layer additional hints over bare ASTs. This paper uses it to add a layer connecting names across an AST, which provides dataflow hints.
To do so, we use the FastGTN model [18], which uses the meta-path, a multi-hop sequence of connected nodes and edges in the graph with heterogeneous edge types. Instead of ignoring the edge type information, breaking down a heterogeneous graph into homogeneous graph by its meta-paths allows the model to preserve this information. A Graph Transformer Network (GTN) [18] learns the most informative meta-paths automatically for each task in an end-to-end manner by selecting adjacency matrices through a 1x1 convolution with weights from a softmax function. The GTN can learn arbitrary meta-paths with respect to edge types and path length. GTNs consider multiple types of meta-paths simultaneously by generating multiple graph structures: they set the output channels of a 1x1 filter to the number of channels. After stacking $k$ GT layers, multi-layer GNNs are applied to each channel of the output tensor. The final node representations can be used for downstream tasks, by applying dense layers followed by a softmax layer to the node representations.
A weakness of GTNs is that they are not scalable because of the high computational cost and memory requirements when calculating and storing new adjacency matrices at each layer. To address this, FastGTNs implicitly transform the graph structures without storing the new adjacency matrices of meta-paths, avoiding the need for explicit multiplications of large adjacency matrices. This reduces the computational cost and memory usage while keeping FastGTNs mathematically identical to GTNs. Even FastGTNs have serious scalability limitations, which motivates our pruning strategies discussed in section III; note that the primary scalability challenge for GTNs is fitting the $n$-by-$n$ adjacency matrix they require into memory (where $n$ is the size of the input)—even with “large” inputs, both training and prediction with FastGTNs is relatively quick (see further discussion in section VI-A3).
II-A3 GCN vs GTN
The GTN is a generalization of a GCN that supports distinct edge types. This allows the GTN to distinguish the AST and the name layer. While GCNs do not automatically support this heterogeneous operation, they are a good baseline for studies exploring advanced graph learning techniques [13]. because they can integrate local neighborhood information through the graph convolution operation.
II-B Text-based Model: the Large Language Model (LLM)
Magicoder [16] is a fully open-source, state-of-the-art LLM trained on a dataset created from 75,000 synthetic instruction data generated using the OSS-INSTRUCT approach. OSS-INSTRUCT involves generating coding problems from real world snippets, ensuring diversity by incorporating a wide variety of challenges, and enhancing the controllability of the output. Since Magicoder has outperformed GPT 3.5 using only open source training samples, it is a suitable baseline for comparison with other techniques.
The LLM approach leverages a vast knowledge base, ensuring a diverse understanding of coding patterns and practices, significantly beneficial for inferring type qualifiers by understanding the broader context and nuances in code. LLMs have recently been used to great effect in a number of software engineering domains [19]. The model’s limitation is that if certain edge cases or less common programming constructs are underrepresented, the model’s inference capability might be constrained for those scenarios. Crucially, LLMs may struggle with deep semantic understanding of code, they are only predicting the next token in a sequence. In practice, there are other limits for using LLMs. For example, Magicoder limits its output to 2048 tokens.
III Encodings
This section details how we encoded the input programs for the two kinds of models we consider: the graph-based models via our novel type-system-specific NaP-AST encoding (which is described via the methodology to construct a NaP-AST encoding for an arbitrary type system in section III-A) and the text-based models via prompting (section III-B).
III-A Graph-based Encoding: the NaP-AST
This section describes our novel methodology for building a type-system-specific program encoding for the graph models (GCN and GTN), which we call a Name-Augmented, Pruned Abstract Syntax Tree (NaP-AST). The encoding only includes information relevant to the given type system, but is in a form suitable for training a deep learning model. The overall methodology for NaP-AST construction is shown in fig. 1.
1. AST Extraction2. Pruning3. Name Augmentation4. Statement Pruning5. Clustering6. Training7. Prediction8. Post-processing
Figure 1: Order of steps in NaP-AST construction: 1. Extract AST (section III-A1). 2. Prune irrelevant nodes or types via proof (section III-A2) and via ablation studies. (section III-A3) 3. Augment the pruned AST by connecting the variables with identical names (section III-A4). 4. Prune irrelevant statements on the basis of relevance (section III-A5). 5. Cluster the ASTs by graph structure (section III-A6). 6/7. Train or predict (section III-A7). 8. If predicting, post-process to ensure consistency (section III-A8).
The base of a NaP-AST is the program’s Abstract Syntax Tree (AST). The first step (section III-A1) to construct a NaP-AST is to augment the AST with additional nodes that model information traditionally stored in the AST directly (for example, whether a declaration is public or private), thereby making that information available to the deep learning model later. This is necessary because the internal complexity of the AST data structure produced by JavaParser is not visible to the deep learning model.
Figure 2: In heterogeneous graphs (i.e. having multiple edge types), a meta-path is a sequence of connected nodes and edges, where each edge has a distinct type.
The next step in constructing a NaP-AST is pruning information that is not relevant to the property of interest from the graph produced by the previous step. We do this by conducting a series of targeted ablation experiments: we iteratively remove particular types of AST nodes (and all associated edges) from the NaP-AST and retrain a preliminary model. Node types whose removal does not negatively impact the F1 score of the preliminary model are removed from all NaP-ASTs. Shrinking the size of the NaP-AST by removing irrelevant information is critical to the scalability of our approach, since the model architectures we consider are highly sensitive to the size of the input graph. There are three pruning phases:
  1. 1.
    prune nodes that are provably unrelated to the property of interest (section III-A2).
  2. 2.
    use an ablation experiment to prune node types that are unrelated to the target property (section III-A3). This pruning step occurs before name augmentation when processing the AST for a particular class.
  3. 3.
    use an ablation experiment on program statement node types to prune entire subtrees under each type of node, as well as related nodes (section III-A5). This pruning step occurs after name augmentation when processing the AST for a particular class, so we defer discussing it until after we introduce the name-augmentation step.
Each pruning phase is theoretically optional, but we found that all three were necessary to produce NaP-AST graphs that are small enough to realistically train a deep learning model when we conducted our feasibility study.
The last part of NaP-AST construction is augmenting the graph with edges that encode program elements that have the same name, which is described in section III-A4. The goal of this step is to encode basic dataflow information in a way that is visible to a deep learning model. For example, this stage enables a model to determine that a parameter checked against null in a method should be annotated as @Nullable by connecting the check to the parameter’s declaration.
Throughout this section, we assume the existence of a training dataset of human-annotated code that contains type qualifiers for the type system of interest. We discuss whether such datasets exist for real type qualifiers in section IV.
III-A1 Abstract Syntax Tree Extraction
For each class in the training dataset, we first generate its Abstract Syntax Tree (AST). Traversing the AST, we generate a graph consisting of nodes and edges. Unlike HiTyper’s Type Dependency Graph [9], we have not attempted to build a graph of type dependencies. We have been able to obtain good results without creating a complex data structure. We think this is because pluggable type systems are much simpler than the type system of a language. Each node is represented by a node id, node type and a special field indicating whether the node contained a type qualifier in the input data (and if, which type qualifier). Any actual AST nodes for the type qualifiers generated by the parser are discarded, so that the human-written type qualifiers are not visible to the model during training. Edges in this graph are parent-child connections represented as an adjacency list.
The graph contains two types of edges: parent-child, and child-parent. Representing the internals of each block or statement by a different edge type prevents relevant properties from leaking out into proximate but otherwise unrelated nodes in the AST. On the other hand, we do want statements using the same variables to be proximate in the graph. We achieve this later using the name layer described in section III-A4.
We then add additional, useful information that is traditionally stored in the AST nodes themselves—in particular, modifiers (e.g., public, static, etc.) —to the graph’s structure, which makes it available to the models during training. We created a single node in the graph for each modifier and connected the modifiers to the relevant nodes using a multi-hot vector encoding. While this may seem like an implementation detail, we mention it here because practical parsers often store this important information not in the AST itself but in fields that are lost if naïvely converted to a graph where each node has one type. We can overcome this issue by representing both the node type and the modifier using multi-hot encoding.
III-A2 Pruning Phase 1: Remove Provably-Unrelated Information
When considering a particular type system, we usually know that some parts of the program are definitely completely irrelevant to the property of interest. If we formalize that notion at the AST level, irrelevant parts of the program can be immediately removed from the AST. For example, a pluggable type system that tracks whether integers are non-negative (which would be useful as e.g., part of a larger system for typechecking array accesses) need not consider parts of the program that do not interact with integers. Therefore, the first pruning stage of NaP-AST construction is to remove any nodes that are provably irrelevant. The set of nodes that are irrelevant will differ when constructing a NaP-AST for each different type system; this pruning stage exists to allow someone following our methodology to build an inference model for a new type system to prune anything they know is irrelevant right away, before the more complex pruning stages.
III-A3 Pruning Phase 2: Ablation by Node Type
Like some Java types are irrelevant for predicting pluggable types, so are some AST node types. For example, MemberValuePair and IntersectionType nodes produced by JavaParser would be irrelevant for predicting non-compound types. To summarize the AST further, we conduct an ablation study on node types. For each node type, we drop it and calculate the F1 score on a held-out test set. We did the same test by dropping random nodes. For node types which get a higher F1 score than dropping random nodes, we drop them from the AST.
III-A4 Name Augmentation
To inform the model of the identity of a candidate node, we layer a graph with different node and edge types on top of the graph derived from the AST in previous steps. This layer connects the names of the variables to nodes that make use of those variables, using different node and edge types. This improves accuracy by allowing the model to connect uses of names to their declarations. Since type annotations need to be placed on declarations, this step is critical: the model must know that a particular test or assignment is related to a specific declaration. For example, fig. 3 shows an example for a variable named "x", which is connected as a separate graph layered over the pruned AST.
The names are connected using a different edge type. The second set of edges are stored in a different adjacency matrix.
This layer can be viewed as encoding a simplistic dataflow analysis, or alternately as enabling the model to reason about basic dataflow. We chose to add this graph layer at the encoding stage rather than running a dataflow analysis directly because it was simpler, and in our experiments it appears to work well in practice. Future work could investigate making dataflow information available to model directly as an alternative to our name augmentation strategy.
MethodDeclarationParameterAssignExpr"x"
Figure 3: Name Augmentation. This diagram shows the pruned AST as yellow nodes, and a node ("x") in the name layer in green. All the nodes that use the same name are connected by a second edge and node type.
III-A5 Pruning Phase 3: Ablation of Statement Subtrees
Even after the pruning phases described above, after name augmentation the resulting graph is prohibitively large to train the graph-based models in section II. So, we need another pruning strategy. Our key insight at this stage is that many statements in the program are irrelevant to the property of interest. For example, parts of the program that do not include any null literals are unlikely to be related to nullability. However, dropping all such statements hurts model performance (as measure by F1 score) too much, so we performed an ablation study to determine which kinds of statements are safe to prune. Our goal at this stage is to prune statements, their corresponding subtrees, and any other related nodes. We use the name-augmentation layer added to the graph in section III-A4 to determine which non-subtree nodes are related to a statement: when we prune a statement and its subtree, we also prune any node that is directly connected to the statement’s subtree via a name layer.
III-A6 Clustering
In most machine learning projects, one cleans the dataset by pruning outliers. Instead of trying to visualize the different kinds of graphs so that we can prune outliers by hand, we clustered the dataset using k-means clustering and trained a separate model for each cluster. We chose a value of k=5 by minimum within-cluster variation with the elbow method [20]. This method raised the overall F1 score by an average of approximately 0.5, so it is critically-important to overall model performance.
III-A7 Prediction
A graph-based model typically makes a prediction for a single input graph (in our case, a NaP-AST representing a single class) at a time. However, program elements defined in one class are sometimes used in other classes in ways that impact their nullability. To capture this phenomenon, we employ a conjoined prediction strategy: for each pair of classes in a target project, we predict the nullability of all of their elements together, using JavaParser’s symbol solver to connect the name layers of the NaP-ASTs if they have any elements in common. The final prediction for an element’s nullability is the average of the predictions across all pairs of classes. This technique increases the consistency of the model’s prediction across classes: conjoined prediction helps when an element is used in more than one class, but makes no difference for elements that are not used in the other class for which predictions are being made concurrently. In other words, this conjoined prediction strategy allows the model to reason (in a limited way) about inter-class dataflow.
III-A8 Post-processing
Deep learning methods are good for finding recommendations through hidden patterns in the data. Unfortunately, their outputs are not guaranteed to be internally consistent. Post-processing steps remove obvious inconsistencies: they certainly do not guarantee that the model’s outputs will be sensible. They should, however, be conservative in a sense: the annotations they add or remove from the model’s output must be guaranteed to make the resulting annotation set for the whole program more consistent.
III-B Text Encoding
To have Magicoder place annotations, we could not provide it with complete source files, because many were too large. We extracted segments of program text that fit in its 2048 token limit (alongside our prompt) using simple rules requiring the resulting text be syntactically complete. Without these rules, Magicoder will break most code trying to fix syntactic errors. Our prompting code’s rules try to include ending braces and complete loops, methods or files where possible.
For each extracted segment, we gave a simple prompt, "Add \$ANNOannotations to the following Java code snippet where having them wouldimprove code quality. Enclose code in ‘‘‘java...‘‘‘" (where \$ANNO is the literal name of the desired annotation: @Nullable in our feasibility study). If "‘‘‘java...‘‘‘" was in the output, the part of the output contained within was written back into the appropriate location in the program. Otherwise, we wrote back the original unannotated code.
This prompting strategy is relatively primitive. There are two reasons for this: first, the prompt is competing for space in the token limit with the code. A longer prompt requires us to provide less code at a time to the model, reducing the model’s performance. Second, the LLM’s performance was poor enough relative to the graph-based models in our preliminary experiments that it did not seem worthwhile to further optimize it. It can therefore be regarded as a baseline in the experiments in the following sections.
IV Dataset
To evaluate the encodings in section III and to answer our other two research questions—i.e., which model is best suited to the task and how much data is actually required—we performed a feasibility study on the most popular pluggable type system: preventing null-pointer dereferences in Java using a @Nullable annotation. Multiple typecheckers for preventing this problem exist and have been deployed in industry, including Uber’s NullAway [2], Meta’s NullSafe [3], and the Checker Framework’s Nullness Checker [4, 5].
The first challenge in predicting nullability annotations is collecting training data. Each nullability checker has its own set of annotations. Further, there are annotations provided by frameworks (e.g., the JSR 305 nullability annotations in the JDK, or those provided by Android) that do not have an associated checking tool. This fragmentation in the ecosystem means that for any single nullability checker or annotation, there may not be sufficient training data to build an effective model. (This mirrors the general situation for pluggable typecheckers, wherein most checkers are used by a relatively small number of projects, leading to a paucity of training data when considering whether training a model is practical.)
Our key insight in dataset preparation is that the various nullability tools all use similar definitions of @Nullable: it universally means “this element may or may not be null.” Since the most widely-used pluggable typecheckers use @NonNull as the default type for unannotated code (e.g., both NullAway [2] and the Checker Framework [4, 5] use this defaulting scheme), bringing a program into compliance with a typechecker usually requires a human to write mostly or only @Nullable annotations. The consequence of this defaulting scheme is that we can restrict a model’s goal to the prediction of only @Nullable annotations: we do not need to directly predict @NonNull or any of the other, more complex annotations used by various checking tools. In turn, the need to only predict @Nullable means that we can combine training data that uses different @Nullable annotations, as long as those annotations have the same semantics. We used every annotation in the comprehensive list of @Nullable annotations with the same semantics in the Checker Framework’s manual [21].
We used the search feature of Sourcegraph’s CLI tool [22]to collect Java classes that import various @Nullable annotations. We searched for @Nullable annotations by themselves. However, this approach led to a dataset where Jetbrains projects were overrepresented. We think this happened because Jetbrains has the most popular projects that use the @Nullable annotation. We switched to searching for a variety of Nullable imports to gather a diverse dataset. We do not require code in our training set be checked by some nullability tool: instead, we try to model where a human would write a nullability annotation. E.g., @Nullable annotations that a human wrote as documentation are also useful as training data.
Some classes, despite importing a nullability annotation, did not actually contain any such annotations. After selecting only those classes that did, we removed very small and very large classes. We removed very small classes because in initial experiments, we found that keeping them in the training set harms model performance: it assumes that all small classes have all @Nullable members. We also removed classes that were too large to fit our training pipeline. After all of these removals, we were left with 32,370 classes; we removed 60,292 classes in total. Each class in this dataset contains one or more @Nullable annotations.
V NaP-ASTs for Nullability
With the dataset in section IV, we trained the graph-based models (GCN and GTN) by following the NaP-AST encoding method described in section III-A. Our code, dataset, and experimental scripts are open-source and available [23]. We call the resulting models NullGCN and NullGTN.
V-A NaP-AST Pruning
In this section, we describe the specific pruning strategies that we used for each pruning phase described in section III-A.
V-A1 Phase 1: Pruning by Proof
Since primitive types cannot store null values in Java, we exclude them from the nodes that the model is trained on. Nullability is relevant to most of the program, but we expect that for other pluggable type systems that target rarer problems, this step could prune dramatically-larger parts of the program. For example, an instantiation of our methodology for a type system that detects resource leaks could remove all parts of the program that definitely do not manipulate resources (i.e., most of the program!).
V-A2 Phase 2: Pruning by Node Type
First, we conducted an ablation study over node types. We dropped each node type and calculated the F1 score of a preliminary model on held-out testing data.The results are plotted in fig. 4 (test runs were repeated 50 times each to control for the statistical nature of the models). Note “<random>” in red in fig. 4: this represents dropping nodes at random from the graph, without regard to their type. We discarded all nodes of the types that appear to the left of “<random>” in fig. 4: that is, node types whose removal damages the model’s predictive ability less than removing random nodes. We connect the children of dropped nodes to their parents.
For example, since humans often write @Nullable annotations on method declaration nodes (to indicate nullable return types), dropping them dramatically reduces the F1 score to about 0.66 . By contrast, node types like MemberValuePair and IntersectionType seem to have no measurable impact on the performance of the model, which makes sense: value pairs are unrelated to reasoning about nullness. As another example, IntersectionType nodes indicate a generic type construct—which is entirely orthogonal to nullness.
Dropping some node types suprisingly drops the F1 score (if ClassOrInterfaceDeclaration is dropped, for example, the F1 score falls to 0.76). We determined via manual investigation that nodes of this type help group the methods of nested classes, so even though they are rarely or never nullable themselves, retaining them is important to model performance. Another node type that helps group nodes together is BlockStmt; dropping them reduces the F1 score to 0.77.
InstanceOfExprArrayInitializerExprTypeExprMethodCallExprForStmtVariableDeclaratorUnaryExprVarTypeNullLiteralExprNormalAnnotationExprIfStmtLambdaExprArrayCreationLevelEnclosedExprBinaryExprImportDeclarationExplCtorInvocStmtReturnStmtTryStmtForEachStmtIntegerLiteralExprWildcardTypeConstructorDeclarationLineCommentConditionalExprObjectCreationExprArrayType<random>ArrayCreationExprArrayAccessExprCharLiteralExprIntersectionTypeAssignExprSingleMemAnnoExprMemberValuePairClassExprParameterEnumDeclarationBlockStmtTypeParameterClassOrInterfaceDeclModifierThisExprMethodDeclaration0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9Node TypeF1 after dropping
Figure 4: Ablation study to decide which node types can be dropped without affecting the F1 score. Only node types that performed worse than dropping random nodes (in red) were retained in later models.
V-A3 Phase 3: Pruning By Subtree
Using a preliminary model, we conducted an ablation study to determine which kinds of statement nodes we can remove, if the corresponding subtree does not contain a NullLiteralExpr node. In the ablation study, we pruned all such subtrees under a particular kind of statement node (and all nodes connected to that subtree via a name layer), for each kind of statement. The resulting F1 scores for the preliminary model appear in fig. 5. Our NaP-AST construction procedure prunes the subtrees and name-connected nodes for the types of statement nodes that appear to the left of the “<random>” entry (in red) in fig. 5: that is, those statement types that are less useful to the model than a random statement. For example, a nullability NaP-AST will not contain a for loop (ForStmt in fig. 5) nor nodes for any of the variables manipulated in that loop unless there is a null literal within the loop. However, all expression statements (not within the subtree of another pruned statement) will be preserved, whether or not the ExpressionStmt node for the statement has a null literal in one of its subtrees—removing expression statements damages model performace too much.
ForStmtBlockStmtLocalClassDeclStmtSynchronizedStmtForEachStmtCatchClauseWhileStmtLabeledStmtAssertStmtThrowStmt<random>ExplCtorInvocStmtReturnStmtIfStmtContinueStmtSwitchStmtTryStmtExpressionStmt0.80.80.80.80.850.850.850.850.90.90.90.90.950.950.950.95Node TypeF1 after dropping
Figure 5: Results of the ablation study described in section III-A5 by statement type. Subtrees of nodes of the types to the left of “<random>”, and nodes connected to those subtrees by the name layer, are pruned from the NaP-AST.
TABLE I: NullAway benchmark results. “War.” means “Warnings.” “Red. %” means “reduction percentage”: that is, the reduction in warnings achieved by NullGTN vs. the unannotated code.
NullGCN: Magicoder: NullGTN:
Project War. Red. % Precision Recall War. Red. % Precision Recall War. Red. % Precision Recall
ColdSnap 2 99% 4 8 Broken NA NA NA 2 99% 15 90
QRContact 0 - 0 0 Broken NA NA NA 0 - 11 61
caffeine 117 -11% 83 44 6 94% 0 0 73 30% 85 95
jib 122 -20% 84 19 Broken NA NA NA 101 0% 62 91
keyvaluestore 44 -450% 13 70 Broken NA NA NA 4 50% 22 100
meal-planner 11 -1000% 0 0 Broken NA NA NA 1 0% 10 100
uLeak 6 50% 37 100 12 0% 0 0 8 33% 75 100
AutoDispose 30 -1400% 3 11 Broken NA 0 0 2 0% 0 0
picasso 0 - 81 44 0 - 25 1 0 - 0 0
RIBs 0 100% 100 100 1 0% 0 0 1 0% 0 0
skaffold-tools-for-java 1 0% 100 83 1 0% 0 0 1 0% 0 0
butterknife 4 0% 100 50 4 0% 100 50 0 100% 100 100
Totals 337 31% .51 .32 NA NA .2 .03 193 60% .6 .89
V-A4 Post-processing
We added four nullability-specific post-processing steps to ensure internal consistency.
Returning Nullable Fields
Our first post-processing step is a simplistic dataflow analysis: when a method returns a field that the model has marked as @Nullable, the method’s return type should be @Nullable, too. This post-processing step is especially important for getters, which the models otherwise rarely predict are nullable.
Consistency between Arguments and Parameters
Our next post-processing step is similar in spirit: at a call-site where an argument is a read of a field that the model predicted is @Nullable, we change the corresponding parameter to @Nullable in the method declaration. We have observed, qualitatively, that these models tend to do a poor job predicting whether parameters are nullable, relative to other kinds of program elements like fields; we conjecture that the reason for this behavior is that nullable fields are often directly assigned null, but many methods with nullable parameters are never directly called with null literal arguments.
Stream Parameters
Since the models use piece-wise boundaries in a continuous, higher dimensional space, their impressionistic understanding of Java syntax occasionally puts annotations where they cause compilation errors. For example, a model might place a @Nullable annotation in a stream parameter, where Java’s syntax does not permit one. We observed this in practice, so we added a post-processing step to remove these @Nullable annotations.
Inheritance
A method’s inferred parameter or return types in the same inheritance hierarchy may not match. We remove @Nullable annotations in superclasses/interfaces no such annotation was inferred in the subclass.
V-B Training
We split the nodes with @Nullable annotations into 60% training, 20% validation, and 20% test datasets. To create a balanced dataset, we had to combine each set with an equal number of nodes without @Nullable annotations. To do this, we took the annotated nodes in the dataset and subjected them to the 60-20-20 split. We combined the training and validation sets with unannotated nodes of equal number from the same classes, and put all remaining nodes in the test set.
TABLE II: GTN hyperparameters
Hyperparameter Value
GT Layers 5
FGTN Layers 2
Epochs 10/batch
Learning Rate 0.01
Channel Agg. Mean
# of Channels 2
Non-local weight -2
k-Hop 9
We trained the GTN on a cluster with Intel Xeon E5-2630-v4 CPUs (2.2 GHz), 2 GPUs and 50 GB RAM. Since the adjacency matrices that we use to represent the AST are sparse, they take up a lot of space on the GPU. So, we were only able to train a sample consisting of 1.15% of our dataset at a time. We thus chose batch training, which allows us to train the GTN on the full dataset over a longer period of time. We fixed the input size to 8000 nodes, and only trained on as many classes at a time as fit into that space. The GCN’s training (and all the experiments in section VI) were run on an AWS g4dn.xlarge instance and a Deep Learning PyTorch AMI.
V-B1 Hyperparameters
After some trial and error, we used a simple 2 layer GCN with batching. The GTN’s trained hyperparameters are in table II. These hyperparameters were chosen by manual grid search. For this dataset, non-local weight of -2 performs well independent of other variables. Since we trained a large dataset in batches, we chose a low initial learning rate and used weight decay throughout the training process to avoid overfitting.
V-C Thresholding during Deployment
The models are trained on an equal number of nodes that are nullable and non-null, and all of its training data was sourced from classes that contained at least one nullable member. Model predictions reflect this: averaged over many pieces of code, the model’s average nullability prediction is about half nullable elements and half non-null elements. In real code, however, there are many more non-null elements than nullable elements: in fact, real code contains many classes that have no nullable members at all! To account for this, when applying each model, we use a thresholding strategy: we only actually mark elements as nullable (by writing an @Nullable annotation) when the model is very confident in that element’s nullability. In our experiments with real projects in section VI-A, we applied a threshold of 90% confidence. In a deployment of one of these models on a previously-unannotated codebase (which is our envisioned deployment scenario), we recommend using a similar threshold.
VI Experiments and Results
We address the research questions from section I:
  • RQ1: What kind of ML model should we use, and how should we post-process its output?
  • RQ2: How should we encode the program in a type-system-specific way to train the model?
  • RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
We performed two experiments. The first is a large-scale evaluation of each model’s ability to infer ground-truth, human-written annotations on real open-source projects from a previous study of a nullability checker (section VI-A) to answer RQ1 and RQ2. The second experiment (section VI-B) evaluated how much training data the best model from section VI-A required for good performance, to answer RQ3.
VI-A NullAway Benchmark
We ran all three models on the open-source benchmark the paper about the NullAway pluggable typechecker [2]. This benchmark (from their artifact [24]) is a collection of 18 projects from GitHub in 2019; each was annotated with nullability annotations when the artifact was published. The artifact also includes an evaluation tool to run NullAway on these projects. Two of these repositories no longer exist;we could build the project and run NullAway on 14 of the remaining 16; two more projects no longer use @Nullable annotations (in the branches pulled by the artifact’s evaluation tool).We consider the remaining 12 projects that we can typecheck using NullAway; however, we note that of these projects, 3 of the 12 do not typecheck with no warnings, even with the human-written annotations. We assume that this is because of either changes in the projects or changes in NullAway itself.
We checked each benchmark project with NullAway in three configurations: 1) with the human-written annotations that came with the projects, 2) after removing the human-written annotations from each project, and 3) after using each model to re-annotate the code. We measured two proxies for human annotation effort: each model’s precision/recall with respect to the human-written annotations (which are erased before passing the projects to each model) and the reduction in the warnings that NullAway issues using our annotations versus the unannotated code. The results appear in table I.
VI-A1 NullGCN
We trained NullGCN using PyTorch and the PyTorch Geometric library. It uses two convolutional layers, dropout regularization, negative log-likelihood loss function and the Adam optimizer. It achieves an accuracy of 0.6 on the test set during training. On the full NullAway benchmark, NullGCN’s performance is not particularly good: its recall is 32%, precision is 51%, and warning reduction is only 31%. While this is better than no inference at all, this result suggests to us that the hypergraph structure of the GTN (which does much better) is important to modeling where humans write annotations: the same graph is not as useful if the edges are not known to be of different types.
VI-A2 Magicoder LLM
We use Magicoder to reannotate the NullAway benchmark projects. Magicoder uses text for input and output, we did not use the NaP-AST encoding.Magicoder seems to perform well on one project (caffeine). However, it often breaks the code in unpredictable ways by making seemingly-nonsensical changes. Even when this does not occur, the LLM’s token limit is too restrictive for it to be effective on most large projects. These problems together lead to its low precision (20%) and recall (3%). While Magicoder appears to do well on the seemingly-difficult caffeine project in table I, the results are difficult to credit: while the resulting code builds with NullAway, and NullAway issues only 6 warnings on it (a seeming warning reduction of 94%!), the structure of the code was altered too much to tell which node goes where: that is, we cannot even compute precision and recall, because Magicoder’s edits to caffeine are too extensive. The LLM did not follow the prompt to not change the code except for placing annotations.
VI-A3 NullGTN
We built our GTN model (NullGTN) by modifying a fork of the original code [25]. It takes an average of about 16 hours to train the model. The largest input source code we evaluated (6,470 non-comment, non-blank lines) took about 5 minutes for inference. NullGTN achieved an F1 score of .95 on the test set during training. On the NullAway Benchmark, NullGTN discovered 89% of the human-written type qualifiers—most of them, with a precision of 60%. However, this only eliminated 60% of the NullAway warnings that a human would need to resolve to adapt these benchmarks to use NullAway—a large amount, but far from all of the warnings. One source of NullGTN’s inability to reduce warnings further is imprecision, which leads to extra annotations, which leads to extra warnings. Another cause of low warning reduction for NullGTN is interactions with libraries. NullGTN has no model of what libraries do, and unlike an LLM it cannot guess based on e.g., the names of library methods. Both the jib and caffeine projects (where NullGTN fails to remove many warnings despite good recall) use libraries extensively: caffeine heavily uses Google Guava, while jib is a modular system with many APIs and plugins, which the model was unaware of.
Overall, these results are encouraging: NullGTN makes a good start on bringing these benchmarks into compliance with NullAway, and finds the vast majority (89%) of the annotations that the developers themselves keep in their codebases. Our answer to RQ2 is therefore positive: NullGTN can make significant progress towards annotating a codebase for a typechecker, and many of the annotations that it produces are remarkably similar to those a human developer would write.
102030405060708090100200200200200400400400400600600600600Training percentageWarnings
Figure 6: Number of warnings by percentage of data trained.
VI-B RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
We rebuilt the GTN model with random samples of the dataset at 10% increments (10%, 20%, …) and ran the evaluation to see how much data is required for good results. The number of warnings by percentage of data trained is shown in fig. 6. What we see is a classic diagram of the model improving with training data until 60%. Performance then worsens after 70% with overfitting. In our feasibility study, performance is maximized from training around 50% of the dataset: around 16,000 classes. This estimate gives us a recommendation to calibrate the dataset size for similar tasks.
VII Limitations and Threats to Validity
VII-A Threats to the Validity of Our Experiments
One threat to our NullAway experiment (section VI-A) is that the quality of the human annotations that we used as ground truth is somewhat questionable: they do not remove all NullAway warnings on 3 of the projects, which may distort our results. Another threat is the use of proxies for human effort: precision/recall of human-written annotations and reduction in warning count may both be poor proxies for effort. For example, even though NullGTN finds most of the annotations a human would write, those it fails to find might be the most difficult or complex to write. Another threat comes from the choice of benchmarks: we use programs that have been, at some point, annotated by a human to typecheck with NullAway, so that we have ground truth for our recall computation. However, in a deployment of NullGTN, we expect that the program would never have been human-annotated. Programs which have never been annotated might be trickier for inference, since human annotators might also change code to simplify the process of annotation (and those changes might stay in the code, even if the annotations themselves are later removed). The use of NullAway itself is also a threat to external validity (generalizability): our approach may not be as effective at inferring annotations that are useful for some other nullness analysis, such as Meta’s NullSafe.
Our approach may not generalize beyond Java and beyond nullability. In theory, our approach should be applicable on other languages beyond Java, but evaluating that is beyond the scope of this work. We expect that the NaP-AST construction procedure in section III-A and a FastGTN architecture will be applicable to other pluggable typecheckers that prove the absence of other sorts of problems in Java, if a dataset of type qualifiers for those checkers were available. However, we cannot test whether this is effective in practice, since as far as we are aware sufficient training data does not exist for typecheckers that target problems other than nullability: there is at least an order of magnitude more code annotated for nullability than for any other typechecker in open-source, and in section VI-B we showed that NullGTN required about half of all the data we collected before it became effective.
VII-B Other Limitations of Our Approach
All machine-learning-based techniques share one serious limitation: there are no guarantees that the annotations that it infers are correct. In our experiments, NullGTN demonstrated close alignment with human-written ground truth annotations—with a recall of 89%. Though there is also no guarantee that human-written annotations are correct, we believe that NullGTN’s ability to hew so closely to developers’ annotations suggests that this problem will be minor in a practical deployment: whether a human or NullGTN writes an incorrect annotation, a sound typechecker will reject it.
VIII Related Work
The most closely-related work to ours is a recent approach for “error-driven” inference of nullability annotations for pluggable typecheckers [26]. The key idea of this appraoch is to treat the typechecker as a “black box” and use its error messages to drive inference: for example, if the checker warns that a return expression is nullable but that a return type is declared as nonnull, this approach could infer that an @Nullable annotation should logically be placed on the return type. In experiments using NullAway, this approach was shown to reduce the number of warnings emitted by the checker by 69.5% (vs. our 60%); they did not measure precision/recall of ground-truth annotations. However, our techniques are different: they relies on the checker’s output, while we rely on where humans have written nullability annotations historically (i.e., the training data for our model). In methodology, the two approaches are orthogonal; we hope to profitably combine them in future work, since the differing approaches mean that they probably infer different subsets of nullability annotations. Another closely related approach bootstraps global inference from the local inference that is built into most pluggable typechecking frameworks [27]. This approach, though not specific to nullness, also uses a fundamentally different approach and might be profitably combined with ours in future work: it works deductively, unlike our approach.
Some prior works [1, 28] use constraint-based type inference for pluggable types. These approaches do not support nullness (they are specific to a units-of-measurement type system), and cannot be readily applied to other pluggable type systems: each constraint-based inference approach is unique. Our methodology, by contrast, could be applied to any type system, if training data were available: a model developer would need to repeat our ablation experiments to drop appropriate nodes from the NaP-AST, but otherwise could reuse our infrastructure. Prior approaches for explaining type inference errors [29, 30, 31, 32] and for gradual type migration [33] suffer the same problems as other constraint-based appraoches (i.e., do not apply easily to new type systems). There is also prior work that involves the user in type inference [34], which is complementary to ours. The Daikon invariant detector [35] can insert @Nullable annotations at program locations that can be null at run time [36]. Like any dynamic approach, it is inherently limited by the quality of program’s test suite.
There is a lot of work on inference of static types for dynamically-typed languages such as Python or JavaScript using machine learning [37, 38, 6, 7, 8, 9, 10, 11]. These approaches are not directly applicable to our problem, since they do not reason about nullability, but about the base type system, and their models and encodings are specialized to that problem.
There is a large literature on type inference and type reconstruction algorithms based on constraint solving [39, Chapter 22]. These techniques build and then solve a system of constraints; the specific constraints are type-system-dependent, which makes this approach unsuitable for pluggable types: determining the constraints to solve would require translating the typechecker’s logic into the logic of the constraint solver, which effectively would require re-implementing the typechecker. Recent work [40] has proposed an extension to the classic constraint-based type inference approach Algorithm W [41] that enables reasoning about nullability. However, extending this to the mainstream object-oriented languages (and their associated nullability checkers) is an open-problem.