### v2

# Nullability Type Inference via Machine Learning

Abstract.
Pluggable type systems improve code safety by adding type qualifiers, but deploying them in legacy Java projects is difficult because annotations must be written by hand. This paper investigates using machine learning to infer type qualifiers. We evaluate graph models—Graph Convolutional Networks (GCNs) and Graph Transformer Networks (GTNs)—and large language models (LLMs). For graphs, we test Type Dependency Graph (TDG) encoding from prior work and introduce a novel encoding, Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), embedding minimal dataflow hints. Models are validated on 9 open-source projects human annotated for the NullAway nullness checker. The GTN with NaP-AST achieves the best downstream impact, cutting checker warnings by 69%, with 69% recall of human-written annotations and 39% precision. The strongest LLM shows the opposite trade-off: 91% precision but only 21% recall, still reducing warnings by 61%. Regarding generalization, we conducted an experiment to estimate the number of classes required in the training set to achieve good performance: performance improves up to roughly 16K annotated classes and then degrades from overfitting around 22K.
pluggable types, machine learning, static analysis
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Conference Name; Month, Year; Locationbooktitle: Proceedings of the Conference Name (Conference Name ’24), Month, Year, Locationisbn: 978-1-4503-XXXX-X/24/08ccs: Software and its engineering
1. Introduction
When programmers write code, they are often aware of semantic properties that a variable must adhere to but which are not checked by the type system of their programming language—for example, that a variable should only hold non-null values. Pluggable type systems (Foster et al., 1999; Bracha, 2004) add custom type qualifiers to a language. A pluggable typechecker statically checks these user-defined qualifiers and warns about potential violations at compile time. For example, PUnits (Xiang et al., 2020) adds qualifiers for scientific units, while Uber’s NullAway (Banerjee et al., 2019b), Meta’s NullSafe (Pianykh et al., 2022), and the Checker Framework’s Nullness Checker (Papi et al., 2008; Dietl et al., 2011) enforce programs do not crash due to null-pointer dereferences with an @Nullable qualifier. A key barrier to bringing legacy codebases into compliance with a pluggable typechecker is that developers must write these qualifiers into their code to see the full benefit. The state of the practice when deploying a new typechecker on a legacy codebase remains manual annotation: that is, type qualifiers are written by hand. The high cost of this manual process discourages developers from using these systems, despite their promise to improve code quality. This paper addresses the problem of bringing a legacy codebase into compliance with a null-safety typechecker by automatically inferring type qualifiers. While our primary focus is on nullability for data availability reasons, we propose a framework that is theoretically adaptable to other finite pluggable type systems, if sufficient training data were available.
Recently, researchers have proposed using machine learning to solve a related, but different, problem: inferring type annotations for Python or JavaScript programs (Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021). This problem shares a key feature with inferring type qualifiers for a legacy codebase: in both cases, the programs were written without feedback from the typechecker, so we expect them not to typecheck cleanly (with any set of annotations). Experimental results suggest that these techniques outperform traditional deductive type inference techniques. Moreover, large language models (LLMs) have recently shown promise on a variety of coding tasks (Roychoudhury et al., 2025; Miserendino et al., 2025; Mündler et al., 2024; Bouzenia et al., 2025; Yang et al., 2024; Ozkaya, 2023; Wu et al., 2023; Xiong et al., 2024; Ross et al., 2023). This paper’s goal is to investigate if similar approaches can work for inferring pluggable types.
However, techniques for inferring types in Python or JavaScript programs are not immediately transferable to our setting, because of three key differences. First, when inferring types for Python or JavaScript, the type system is fixed (it enforces “traditional” type safety: objects do not try to perform invalid operations). A key advantage of pluggable type systems is their customizability: each type system enforces its own rules. So, we desire a way to customize our inference approach that does not require us to start over from scratch for each new type system. Second, when inferring types for e.g., Python, the space of possible types is infinite and has a long tail, because of the presence of user-defined types. Prior works focus on solving this so-called rare type problem; for example, by encoding programs using a Type Dependency Graph (TDG(Peng et al., 2022b) data structure which is specialized to help the model find rare types. However, while the tail of types is long, the set of locations that need a type annotation is not: in prior work, the locations where a type should be written are fixed in advance. For most pluggable type systems, these problems are reversed: the hardest problem is not choosing which type to write, but at which places to write a type. For example, for a nullability type system, an annotator usually only needs to write either @Nullable or @NonNull (and the pluggable type system interprets the lack of a qualifier as the other), so the problem that an inference tool needs to solve is where to write a qualifier rather than which type to write. This difference makes the encodings from prior works (e.g., the TDG) ill-suited to our setting. Third, machine learning is data hungry: approaches for popular languages like Python can assume the existence of a large corpus of human-written types in open-source software. The customizability of pluggable types is a double-edged sword: many pluggable type systems are highly-specialized and therefore have relatively few users. So, to apply machine learning techniques to the pluggable type inference setting, we need to understand how much data is actually required to make them effective. These differences inspired us to investigate two high-level research questions:
  1. RQ1
    What kind of machine-learning model architecture is suitable for inferring pluggable types, and how does that model differ from those in prior type inference works for other languages?
  2. RQ2
    How much data is needed to make that model effective at inferring qualifiers in practice?
The results of our investigation of these questions allows us to propose a general pipeline for training an ML model to infer type qualifiers for a given pluggable type system.
State-of-the-art models (Peng et al., 2023b; Wei et al., 2023) for inferring types in Python/JavaScript use diverse model architectures, including graph and text-based models: Typilus (Allamanis et al., 2020) learns a type space, STIR (Peng et al., 2023b) uses a BiLSTM for incomplete code, and TypeT5 (Wei et al., 2023) uses a pre-trained seq2seq model. The differences with the pluggable types setting, however, inspired us to investigate RQ1 by evaluating a set of models used in previous work. We eschew time-series models like BiLSTM (which are effective for sequential data, like recognizing the content of a video, but very expensive during training and inference) because predicting where to write a type qualifier is more similar to a classification task. We considered four models: two graph models and two text-based models. For the graph models, we considered two encodings: the TDG encoding (Peng et al., 2022b) from prior work and a novel encoding that we developed specifically for pluggable type inference. Our novel graph encoding methodology uses ablation studies to retain information useful for predicting the type qualifiers of a particular type system, which results in a novel program encoding that we call a Name-augmented, Pruned Abstract Syntax Tree (NaP-AST), which we describe further in section 3. A NaP-AST encodes the parts of the program’s structure that are relevant to the type system of interest, along with name information that ties related program elements together, into a graph that is a suitable input to a graph-based model. The first graph model is a baseline (Wu et al., 2019) graph convolutional network (GCN) model (Zhang et al., 2019a). The second is a more-capable graph transformer network (GTN) model (Yun et al., 2019), which extends the GCN to reason about multiple edge types (i.e., hypergraphs). Since the TDG encoding is not a heterogenous graph, we only use it with the GCN model; the NaP-AST encoding is a heterogenous graph, so it can benefit from the GTN model. Both text-based models are large language models (LLMs), which we include because LLMs have performed well on coding tasks recently. The first is CodeQwen1.5-7B-Chat  [^1] https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat, an open-source model specialized for coding tasks. The second text-based model is GPT-4 (Achiam et al., 2023), a state-of-the-art [^2] GPT-4 was OpenAI’s flagship model when these experiments were conducted, though at the time of submission OpenAI recently (August 7, 2025) released its new GPT-5 model. commercial LLM.
To evaluate these models and encodings, we conducted a feasibility study: we trained models to predict where to place type qualifiers for one popular pluggable type system. We would have preferred to repeat this experiment for more than one type system, to show the generality of our approach, but a lack of publicly-available data prevents us from doing so; we show this empiricially in section 7.3.1. Of the commonly used type qualifiers, those for nullability in Java are the most common. We therefore conducted our feasibility study on nullness: if there is not enough data for Java nullness type qualifiers to train an effective model, there is not enough data to train an effective model for any pluggable type system. First, we collected a dataset from GitHub of Java classes, each of which uses one of 31 different @Nullable annotations. After cleaning, this dataset contains 32,370 classes. We then used our methodology for building a NaP-AST encoding to derive a NaP-AST encoding for nullness. Then, we trained the four models using this dataset: the two graph models (a GCN and a GTN) using the derived NaP-AST encoding, the GCN using the baseline TDG encoding (which cannot benefit from the GTN, because it is a homogenous graph), and the text-based LLMs (CodeQwen1.5-7B-Chat and GPT-4) using syntax-based code chunking. We then compared their performance on 9 open-source projects that had previously been annotated for the popular NullAway typechecker (Banerjee et al., 2019b). The GTN with the NaP-AST encoding outperformed the other models we evaluated: both GCN models have lower recall than NullGTN; TDG GCN has higher precision than NullGTN but fails to reduce warnings, while NullGCN trails on both precision and recall. LLMs can hallucinate edits at scale; CodeQwen attains very high precision (0.91) but low recall (0.21), and GPT-4 has high precision (0.76) with moderate recall (0.25). Our best model (which we call NullGTN) achieves 69% recall of the human-written type qualifiers and 39% precision; its inferred annotations eliminate 69% of the NullAway warnings that a human would need to triage (a proxy for human effort).
Our success for nullability suggests that our approach may be effective with other type qualifiers, but unfortunately we lack enough data for any other type qualifier to test this hypothesis. To assist future researchers or practitioners who want to train a similar model for another type system (and to answer our RQ2), we explore how much training data one needs for adequate performance. We rebuilt our best model with random samples of the dataset and re-ran the evaluation to show how much data it needs before it achieves good performance. Around 16k human-annotated classes were needed: a large but not impossible-to-reach number. Our primary contributions are:
  • A novel methodology to construct a graph encoding of programs for null safety (NaP-ASTs) to train deep-learning graph models like GCNs and GTNs that can theoretically be extended to other finite type systems (section 3);
  • A dataset of 32,370 Java classes annotated with @Nullable that enables our feasibility study (section 4);
  • A comparison of three model architectures (GCN, GTN, and LLM) and two graph encodings (NaP-AST and TDG) that shows that the GTN with the NaP-AST encoding is best at inferring pluggable types in our feasibility study (section 7.1), with open-source models and data;
  • A study of the amount of data required by our best GTN model on the task of inferring type qualifiers for adequate performance (section 7.3).
2. Background: Model Architectures
This section describes the three considered model architectures: graph convolutional networks (GCNs), graph transformer networks (GTNs), and large language models (LLMs). We focus on their salient features with respect to type qualifier inference.
2.1. Graph Models
2.1.1. GCN
Graph Convolutional Networks (GCNs) are a class of neural networks for graph data. Type qualifier prediction can be modeled as a graph data task since ASTs can be represented as a graph. GCNs use graph convolution, a propagation rule that updates the feature representation of each node by aggregating features from its neighbors and the node itself. For AST prediction, GCNs can be particularly effective, since they explicitly model local context: ASTs are inherently hierarchical and structured, where the context around a node (e.g., its parents or children) significantly influences it. GCNs leverage this local structure efficiently. They are also efficient in learning spatial features, which have been helpful for ASTs (Zhang et al., 2019b).
A more specialized graph neural network model might outperform a GCN by adapting to heterogeneous relationships. If different types of connections between AST nodes (e.g., data flow vs. control flow) play a significant role, models within the broader GNN category designed to handle such heterogeneity might be more effective. We investigate this possibility by comparing GCN with such a model, GTN. Certain GNN architectures might also be better suited for incorporating global context or long-range dependencies within the AST, which could encode dataflow information.
2.1.2. GTN
GCNs assume their input graphs are homogeneous, i.e., they consist of one type of nodes and edges. Using a model that supports heterogenous graphs lets us layer hints over the AST, which we use to add a layer connecting names across an AST to provide dataflow hints.
We use the FastGTN model (Yun et al., 2022), which uses the meta-path, a multi-hop sequence of connected nodes and edges in the graph with heterogeneous edge types. Instead of ignoring the edge type information, breaking down a heterogeneous graph into homogeneous graph by its meta-paths allows the model to preserve this information. A Graph Transformer Network (GTN) (Yun et al., 2022) learns the most informative meta-paths automatically for each task in an end-to-end manner by selecting adjacency matrices through a 1x1 convolution with weights from a softmax function. The GTN can learn arbitrary meta-paths with respect to edge types and path length. GTNs consider multiple types of meta-paths simultaneously by generating multiple graph structures: they set the output channels of a 1x1 filter to the number of channels. After stacking $k$ GT layers, multi-layer GNNs are applied to each channel of the output tensor. The final node representations can be used for downstream tasks, by applying dense layers followed by a softmax layer to the node representations.
MethodDeclSimpleName”f”SimpleNameMethodCallExprAST ChildName ParentName ChildAST Parent
Figure 1. In heterogeneous graphs (i.e. having multiple edge types), a meta-path is a sequence of connected nodes and edges, where each edge has a distinct type.
A weakness of GTNs is that they are not scalable because of the high computational cost and memory requirements when calculating and storing new adjacency matrices at each layer. To address this, FastGTNs implicitly transform the graph structures without storing the new adjacency matrices of meta-paths, avoiding the need for explicit multiplications of large adjacency matrices. This reduces the computational cost and memory usage while keeping FastGTNs mathematically identical to GTNs. Even FastGTNs have serious scalability limitations, which motivates our pruning strategies discussed in section 3; note that the primary scalability challenge for GTNs is fitting the $n$-by-$n$ adjacency matrix they require into memory (where $n$ is the size of the input)—even with “large” inputs, both training and prediction with FastGTNs is relatively quick (see further discussion in section 7.2).
2.1.3. GCN vs GTN
The GTN is a generalization of a GCN that supports distinct edge types. This allows the GTN to distinguish the AST and the name layer. While GCNs do not automatically support this heterogeneous operation, they are a good baseline for studies exploring advanced graph learning techniques (Wu et al., 2019) because they can integrate local neighborhood information through the graph convolution operation.
2.1.4. Models from Prior Work
Although we did test the TDG encoding (section 3.1) from prior work, we did not evaluate the Python-specific models with which it was originally deployed, because they are not transferable to our setting. In particular, the Typilus paper (Allamanis et al., 2020) used a custom “TypeSpace”, metric-based meta-learning, and a Graph Neural Network (GNN) model optimized for predicting rare types. None of this machinery is sensible or necessary for predicting pluggable type qualifiers, which are typically drawn from a fixed set—for example, in our feasibility study, from the set containing @Nullable and @NonNull.
2.2. Text-based Models
2.2.1. CodeQwen1.5-7B-Chat
CodeQwen 1.5-7B-Chat [^3] https://qwenlm.github.io/blog/codeqwen1.5/ is a fully open-source transformer (7 B params, 64 K context) trained on 3 T code tokens spanning 92 programming languages, and attains 83.5% on HumanEval-plus and 78.7% on MBPP-plus, outscoring other 7B code models  [^4] https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat. The model is fine-tuned with a synthetic instruction corpus that couples long-context code snippets with conversational tasks (Bai et al., 2023). Because CodeQwen 1.5-7B-Chat outperforms StarCoder 2, CodeLlama 7 B and DeepSeek-Coder on major code benchmarks, it is an appropriate open-source baseline. The LLM approach leverages a vast knowledge base, ensuring a diverse understanding of coding patterns and practices, significantly beneficial for inferring type qualifiers by understanding the broader context and nuances in code. LLMs have recently been used to great effect in a number of software engineering domains (Roychoudhury et al., 2025; Miserendino et al., 2025; Mündler et al., 2024; Bouzenia et al., 2025; Yang et al., 2024; Ozkaya, 2023; Wu et al., 2023; Xiong et al., 2024; Ross et al., 2023). Their limitation is that they may struggle with deep semantic understanding of code, as they are only predicting the next token in a sequence. Specifically, LLMs hallucinate about code structure. These hallucinations add up when scaled to hundreds of thousands of lines of code. While CodeQwen benefits from the 64 K-token window, it can still hallucinate nonsensical changes, like adding or removing import statements in large files. All current LLM systems hallucinate since their function is to predict the next token. To get around this limitation, we only asked the LLM “yes or no” questions, restricting its ability to change code (see section 3.3 for more details). In practice, there are other limits on LLMs, like token limits and cost.
2.2.2. GPT-4o-mini
GPT-4o-mini is an LLM that demonstrates exceptional performance in coding and natural language processing tasks, outperforming its predecessors and many contemporary models in various domains. For example, GPT-4o-mini achieves superior accuracy in programming competitions and exhibits strong cross-language translation capabilities, showcasing its adaptability and generalization skills (Hou and Ji, 2024). Its ability to refactor and annotate code with significant quality improvements further highlights its utility in software engineering (Poldrack et al., 2023).
Although excelling in generalization, GPT-4o-mini still faces notable limitations. Its reliance on token limits restricts its ability to process large datasets in a comprehensive way, and exhibits a tendency to hallucinate or produce errors in domain-specific contexts, such as radiology or highly technical programming tasks (Liu et al., 2023; Pordanesh and Tan, 2024). Like with CodeQwen, we restricted its ability to change code, asking only yes or no questions.
3. Encodings
serialize()bytes (byte[])has_parameter
Figure 2. An example TDG snippet. A function $serialize()$ that has a parameter called bytes of type $byte[]$. $serialize()$ is an expression node. bytesbytes is a symbol node. $Has_{p}arameter$ is a merge node.
This section details the three encodings that we considered. The first two encodings are for the graph model architectures (GCNs and GTNs); the last encoding is for the text-based LLM models:
  • the type dependency graph (TDG) encoding from prior work (section 3.1),
  • our novel type-system-specific NaP-AST encoding (which is described via the methodology to construct a NaP-AST encoding for an arbitrary type system in section 3.2), and
  • the text-based encoding via prompting (section 3.3).
3.1. Type Dependency Graph
Prior work (Peng et al., 2022b) used the TDG encoding for inferring type annotations for programs in dynamically-typed languages such as Python. The TDG is our baseline encoding for graph models: it has worked for a similar problem in the past. A TDG has four node types: 1) symbol nodes, which represent variables for which the types need to be inferred. Each occurrence of a variable is treated as a unique node to handle scenarios where the variable’s type may change at runtime. 2) expression nodes, which represent expressions that generate types, such as arithmetic operations, function calls, and list comprehensions. 3) branch nodes, which represent branching points in data flow, typically corresponding to conditional statements. 4) merge nodes, which represent merging points in data flow, such as the convergence of different branches of execution. Our implementation of TDGs for Java involved specific adjustments to handle Java-specific syntax and semantics. For example, we have removed the possibility that a variable’s types may change at runtime. We trained TDGs on a GCN model. A GTN would not confer much benefit since the edges of a TDG have relatively homogeneous semantics. However, the number of types inferred for a pluggable type system is small in comparison to the use case a TDG was designed for. Therefore, the main functionality proposed by these systems is underutilized for predicting pluggable types, motivating our NaP-AST encoding. An example TDG is shown in fig. 2.
3.2. Our Novel Graph Encoding: the NaP-AST
This section describes our novel methodology for building a type-system-specific program encoding for the graph models (GCN and GTN), which we call a Name-Augmented, Pruned Abstract Syntax Tree (NaP-AST). Unlike a TDG, this encoding only includes information relevant to the property being enforced by a particular pluggable type system, in a form suitable for training a deep-learning model. The overall methodology for NaP-AST construction is shown in fig. 3. We explain how we built a NaP-AST encoding for a nullability type system in detail later, in section 5; this section gives an overview of the process of NaP-AST construction for an arbitrary pluggable type system.
The base of a NaP-AST is the program’s Abstract Syntax Tree (AST). The first step (section 3.2.1) to construct a NaP-AST is to augment the AST with additional nodes that model information traditionally stored in the AST directly (for example, whether a declaration is public or private), thereby making that information available to the deep-learning model later. This is necessary because the internal complexity of the AST data structure produced by JavaParser is not visible to the deep-learning model.
1. AST Extraction2. Pruning3. Name Augmentation4. Statement Pruning5. Clustering6. Training7. Prediction8. Post-processing
Figure 3. Order of steps in NaP-AST construction.
The next step in constructing a NaP-AST is pruning information that is not relevant to the property of interest from the graph produced by the previous step. We do this by conducting a series of targeted ablation experiments: we iteratively remove particular types of AST nodes (and all associated edges) from the NaP-AST and retrain a preliminary model using a small dataset (976 classes, in our instantiation of this process; this number is related to batch sizes and memory limitations on the training machine).
The F1 score was calculated by comparing the predicted nodes with the ground truth using the values in the confusion matrix, namely the harmonic mean of precision and recall. The F1 score on this preliminary model was used to determine which nodes to prune.
Node types whose removal does not negatively impact the F1 score of the preliminary model are removed from all NaP-ASTs. Shrinking the size of the NaP-AST by removing irrelevant information is critical to the scalability of our approach, since the model architectures we consider are highly sensitive to the size of the input graph. There are three pruning phases:
  1. (1)
    prune nodes that are provably unrelated to the property of interest (section 3.2.2).
  2. (2)
    use an ablation experiment to prune node types that are unrelated to the target property (section 3.2.3). This pruning step occurs before name augmentation when processing the AST for a particular class.
  3. (3)
    use an ablation experiment on program statement node types to prune entire subtrees under each type of node, as well as related nodes (section 3.2.5). This pruning step occurs after name augmentation when processing the AST for a particular class, so we defer discussing it until after we introduce the name-augmentation step.
Each pruning phase is theoretically optional, but we found that all three were necessary to produce NaP-AST graphs that are small enough to realistically train a deep-learning model. If the pruning steps were not applied, less relevant code would be available to the model in each step (because the model is memory-bound). This makes the model’s performance suffer during prediction: in our early experiments on a subset of the full dataset, we found that omitting pruning reduced the model’s F1 score from 0.95 to 0.78.
The last NaP-AST construction step is augmenting the graph with edges encoding program elements with the same name, which is described in section 3.2.4. The goal of this step is to encode basic dataflow information in a way that is visible to a deep-learning model. For example, this stage lets a model determine that a parameter checked against null in a method should be annotated as @Nullable by connecting the check to the parameter’s declaration.
In this section, we assume the existence of a training dataset of human-annotated code that contains type qualifiers for the type system of interest. We discuss whether such datasets exist for real type qualifiers in section 4.
3.2.1. Abstract Syntax Tree Extraction
For each class in the training dataset, we first generate its Abstract Syntax Tree (AST). Traversing the AST, we generate a graph consisting of nodes and edges. Unlike HiTyper’s Type Dependency Graph (Peng et al., 2022a), we have not attempted to build a graph of type dependencies. We have been able to obtain good results without creating a complex data structure. We think this is because pluggable type systems are much simpler than the type system of a language. Each node is represented by a node id, node type and a special field indicating whether the node contained a type qualifier in the input data (and if, which type qualifier). Any actual AST nodes for the type qualifiers generated by the parser are discarded, so that the human-written type qualifiers are not visible to the model during training. Edges in this graph are parent-child connections represented as an adjacency list.
The graph contains two types of edges: parent-child, and child-parent. Representing the internals of each block or statement by a different edge type prevents relevant properties from leaking out into proximate but otherwise unrelated nodes in the AST. On the other hand, we do want statements using the same variables to be proximate in the graph. We achieve this later using the name layer described in section 3.2.4.
We then add additional, useful information that is traditionally stored in the AST nodes themselves—in particular, modifiers (e.g., public, static, etc.) —to the graph’s structure, which makes it available to the models during training. While this may seem like an implementation detail, we mention it here because practical parsers often store this important information not in the AST itself but in fields that remain unrepresented in the graph structure. We represent this information by creating a single node in the graph for each modifier and connecting the modifiers to the relevant nodes using a multi-hot vector encoding.
3.2.2. Pruning Phase 1: Remove Provably-Unrelated Information
When considering a particular type system, we usually know that some parts of the program are definitely completely irrelevant to the property of interest. If we formalize that notion at the AST level, irrelevant parts of the program can be immediately removed from the AST. For example, a pluggable type system that tracks whether integers are non-negative (which would be useful as e.g., part of a larger system for typechecking array accesses) need not consider parts of the program that do not interact with integers. Therefore, the first pruning stage of NaP-AST construction is to remove any nodes that are provably irrelevant. For example, primitive types cannot store null values in Java, and so NullGTN removes them at this stage. The set of nodes that are irrelevant will differ when constructing a NaP-AST for each different type system; this pruning stage exists to allow someone following our methodology to build an inference model for a new type system to prune anything they know is irrelevant right away, before the more complex pruning stages.
3.2.3. Pruning Phase 2: Ablation by Node Type
Like some Java types are irrelevant for predicting pluggable types, so are some AST node types. For example, MemberValuePair and IntersectionType nodes are irrelevant for predicting non-compound types. To summarize the AST further, we conduct an ablation study on node types. For each node type, we drop it and calculate the F1 score on a held-out test set. We did the same test by dropping random nodes. We drop node types which score worse than dropping random nodes from the AST. Note that when we prune a node, we connect the node’s children to its parents.
3.2.4. Name Augmentation
To inform the model of the identity of a candidate node, we layer a graph with different node and edge types on top of the graph derived from the AST in previous steps. This layer connects the names of the variables to nodes that make use of those variables, using different node and edge types. This improves accuracy by allowing the model to connect uses of names to their declarations. Since type annotations need to be placed on declarations, this step is critical: the model must know that a particular test or assignment is related to a specific declaration. For example, fig. 4 shows an example for a variable named ”x”, which is connected as a separate graph layered over the pruned AST. The names are connected using a different edge type. The second set of edges are stored in a different adjacency matrix.
This layer can be viewed as encoding a simplistic dataflow analysis, or alternately as enabling the model to reason about basic dataflow. We chose to add this graph layer at the encoding stage rather than running a dataflow analysis directly because it was simpler, and in our experiments it appears to work well in practice. Future work could investigate making dataflow information available to model directly as an alternative to our name augmentation strategy.
3.2.5. Pruning Phase 3: Ablation of Statement Subtrees
Even after the pruning phases described above, after name augmentation the resulting graph is prohibitively large to train the graph-based models in section 2. So, we need another pruning strategy. Our key insight at this stage is that many statements in the program are irrelevant to the property of interest. For example, parts of the program that do not include any null literals are unlikely to be related to nullability. However, dropping all such statements hurts model performance (as measure by F1 score) too much, so we performed an ablation study to determine which kinds of statements are safe to prune. Our goal at this stage is to prune statements, their corresponding subtrees, and any other related nodes. We use the name-augmentation layer added to the graph in section 3.2.4 to determine which non-subtree nodes are related to a statement: when we prune a statement and its subtree, we also prune any node that is directly connected to the statement’s subtree via a name layer.
3.2.6. Clustering
In machine learning projects, one typically cleans the dataset by pruning outliers. Instead of trying to visualize the different kinds of graphs so that we can prune outliers by hand, we clustered the dataset using k-means clustering and trained a separate model for each cluster. We chose a value of k=5 by minimum within-cluster variation with the elbow method (Cui et al., 2020). This method raised the overall F1 score by approximately 0.5, so it is critically-important to overall model performance.
3.2.7. Prediction
A graph-based model typically makes a prediction for a single input graph (in our case, a NaP-AST representing a single class) at a time. However, program elements defined in one class are sometimes used in other classes in ways that impact their nullability. To capture this phenomenon, we employ a conjoined prediction strategy: for each pair of classes in a target project, we predict the nullability of all of their elements together, using JavaParser’s symbol solver to connect the name layers of the NaP-ASTs if they have any elements in common. The final prediction for an element’s nullability is the average of the predictions across all pairs of classes. This technique increases the consistency of the model’s prediction across classes: conjoined prediction helps when an element is used in more than one class, but makes no difference for elements that are not used in the other class for which predictions are being made concurrently. In other words, this conjoined prediction strategy allows the model to reason (in a limited way) about inter-class dataflow.
MethodDeclarationParameterAssignExpr”x”
Figure 4. Name Augmentation. This diagram shows the pruned AST as yellow nodes, and a node (”x”) in the name layer in green. All the nodes that use the same name are connected by a second edge and node type.
3.2.8. Post-processing
Deep-learning methods are good for finding recommendations through hidden patterns in the data. Unfortunately, their outputs are not guaranteed to be internally consistent. Post-processing steps remove obvious inconsistencies: they certainly do not guarantee that the model’s outputs will be sensible. They should, however, be conservative in a sense: the annotations they add or remove from the model’s output must be guaranteed to make the resulting annotation set for the whole program more consistent.
3.2.9. Other types
The NaP-AST encoding can be adapted with relevant changes to any type system with finite types. For example, consider a type system for the signedness of integers. The relevant node types, etc., would almost certainly differ from the things that matter for null safety (e.g., null literals might not matter, but addition or subtraction nodes might matter a lot), so the process that we describe in section 3 would need to be repeated. We separated the description of the process in section 3 from the instantiation of that process for null safety in section 5 to facilitate repeating our approach in the future.
3.3. Text Encoding
To have LLMs place annotations, we could not provide it with complete source files, because many were too large. We extracted segments that fit within a conservative  2 K-token budget to keep inference consistent across models and reduce nonsensical outputs, even though some models (e.g., CodeQwen 1.5-7B-Chat) support much larger contexts (up to 64 K).
To avoid hallucinations that change code structure, we restrict the output of LLMs to yes or no, using them as classifiers. First, the script extracts potential candidates for @Nullable annotation from the Java file by scanning its lines with two regular expressions, field declarations and method signatures. The latter captures both return types and parameter lists. For each match, a unique identifier and their corresponding elements are stored in a dictionary for further processing. Next, a prompt is constructed consisting of a concise system message (”You are a helpful assistant.”) followed by a detailed user message. The user message provides the full Java source code and lists each extracted candidate element, prefixed by its unique key. The LLM is then asked to decide for each candidate whether it should be annotated with @Nullable. The script explicitly instructs the LLM to return a strictly formatted JSON object that maps each candidate key to either ”yes” or ”no.” For any candidate marked with ”yes,” it inserts the @Nullable annotation in the appropriate location, whether at a field, method return, or parameter. Without these guardrails, we found that the LLMs break most code segments trying to make syntactic changes—for example, by adding or removing import statements unnecessarily. Our prompting code’s rules try to include ending braces and complete loops, methods or files where possible to help reduce the impact of these problems.
4. Dataset
To evaluate the encodings in section 3 and to answer our other two research questions—i.e., which model is best suited to the task and how much data is actually required—we performed a feasibility study on the most popular pluggable type system: preventing null-pointer dereferences in Java using a @Nullable annotation. Multiple typecheckers for preventing this problem exist and have been deployed in industry, including Uber’s NullAway (Banerjee et al., 2019b), Meta’s NullSafe (Pianykh et al., 2022), and the Checker Framework’s Nullness Checker (Papi et al., 2008; Dietl et al., 2011).
The first challenge in predicting nullability annotations is collecting training data. Each nullability checker has its own set of annotations. Further, there are annotations provided by frameworks (e.g., the JSR 305 nullability annotations in the JDK, or those from Android) without an associated checking tool. This fragmentation in the ecosystem means that for any single nullability checker or annotation, there is not enough training data to build an effective model. This mirrors the general situation for pluggable typecheckers, wherein most checkers are used by only a small number of projects, leading to too little training data when considering whether training a model is practical.
Our key insight in dataset preparation is that the various nullability tools all use similar definitions of @Nullable: it universally means “this element may or may not be null.” Since the most widely-used pluggable typecheckers use @NonNull as the default type for unannotated code (e.g., both NullAway (Banerjee et al., 2019b) and the Checker Framework (Papi et al., 2008; Dietl et al., 2011) use this defaulting scheme), bringing a program into compliance with a typechecker usually requires a human to write mostly or only @Nullable annotations. Consequently, we can restrict a model’s goal to the prediction of only @Nullable annotations: we do not need to directly predict @NonNull or any other, more complex annotations. In turn, the need to only predict @Nullable means that we can combine training data that uses different @Nullable annotations, as long as those annotations have the same semantics. We used every annotation in the comprehensive list of @Nullable annotations with the same semantics in the Checker Framework’s manual (cf-, 2024).
We used the search feature of Sourcegraph’s CLI tool (Slack and Liu, 2013) to collect Java classes that import @Nullable annotations. We searched for @Nullable annotations by themselves. However, this approach led to a dataset where JetBrains projects were overrepresented. We think this happened because Jetbrains has the most popular projects that use the @Nullable annotation. We switched to searching for a variety of Nullable imports to gather a diverse dataset. We do not require code in our training set be checked by some nullability tool: instead, we try to model where a human would write a nullability annotation. E.g., @Nullable annotations that a human wrote as documentation are also useful as training data.
Some classes, despite importing a nullability annotation, did not actually contain any such annotations. After selecting only those classes that did, we removed very small and very large classes. We define “very small” classes as those with fewer than 300 AST nodes. There are 11,019 such classes in the dataset. Based on manual review, we observed that these classes lacked nontrivial logic (besides getters, setters and annotated fields). Our early experiments showed that retaining them in the training set skews the model towards annotating every annotatable location. We define “very large” classes as those with more than 8,000 nodes after pruning. We decided to remove these because we wanted the model to see the complete code of each class during training. 8,000 nodes was the empirical limit imposed by the memory capacity of the training machine. After all of these removals, we were left with 32,370 classes; we removed 60,292 classes in total. Note that despite excluding both categories during training, the models operate on all classes regardless of size when deployed. The NullAway benchmark (section 7.1) contains 1641 “too small” classes and 24 “too large” classes that would have been removed during training. Our results there show that removing these classes during training does not cause problems during deployment. Each class in this dataset contains one or more @Nullable annotations. There are 217,922 @Nullable annotations in these classes in total.
5. NaP-ASTs for Nullability
With the dataset in section 4, we trained graph-based models (GCN and GTN) using our NaP-AST encoding method (section 3.2). This section describes the NaP-AST construction process; we call the resulting models NullGCN and NullGTN.
5.1. NaP-AST Pruning
In this section, we describe the specific pruning strategies that we used for each pruning phase described in section 3.2.
5.1.1. Phase 1: Pruning by Proof
Since primitive types cannot store null values in Java, we exclude them from the nodes that the model is trained on. Nullability is relevant to most of the program, but we expect that for other pluggable type systems that target rarer problems, this step could prune dramatically-larger parts of the program. For example, an instantiation of our methodology for a type system that detects resource leaks could remove all parts of the program that definitely do not manipulate resources (i.e., most of the program).
5.1.2. Phase 2: Pruning by Node Type
InstanceOfExprArrayInitializerExprTypeExprMethodCallExprForStmtVariableDeclaratorUnaryExprVarTypeNullLiteralExprNormalAnnotationExprIfStmtLambdaExprArrayCreationLevelEnclosedExprBinaryExprImportDeclarationExplCtorInvocStmtReturnStmtTryStmtForEachStmtIntegerLiteralExprWildcardTypeConstructorDeclarationLineCommentConditionalExprObjectCreationExprArrayType¡random¿ArrayCreationExprArrayAccessExprCharLiteralExprIntersectionTypeAssignExprSingleMemAnnoExprMemberValuePairClassExprParameterEnumDeclarationBlockStmtTypeParameterClassOrInterfaceDeclModifierThisExprMethodDeclaration0.60.60.70.70.80.80.90.9Node TypeF1 after dropping
Figure 5. Ablation study to decide which node types can be dropped without affecting the F1 score. Only node types that performed worse than dropping random nodes (in red) were retained in later models.
First, we conducted an ablation study over node types. We dropped each node type and calculated the F1 score of a preliminary model on held-out testing data.The results are plotted in fig. 5 (test runs were repeated 50 times each to control for the statistical nature of the models). Note “¡random¿” in red in fig. 5: this represents dropping nodes at random from the graph, without regard to their type. We discarded all nodes of the types that appear to the left of “¡random¿” in fig. 5: that is, node types whose removal damages the model’s predictive ability less than removing random nodes. We connect the children of dropped nodes to their parents.
For example, since humans often write @Nullable annotations on method declarations (to indicate nullable return types), dropping them dramatically reduces the F1 score to about 0.66 . By contrast, node types like MemberValuePair and IntersectionType seem to have no measurable impact on the performance of the model, which makes sense: value pairs are unrelated to reasoning about nullness. As another example, IntersectionType nodes indicate a generic type construct—which is entirely orthogonal to nullness.
Dropping some node types surprisingly drops the F1 score (if ClassOrInterfaceDeclaration is dropped, for example, the F1 score falls to 0.76). We determined via manual investigation that nodes of this type help group the methods of nested classes, so even though they are rarely or never nullable themselves, retaining them is important to model performance. Another node type that helps group nodes together is BlockStmt; dropping them reduces the F1 score to 0.77.
5.1.3. Phase 3: Pruning By Subtree
Using a preliminary model, we conducted an ablation study to determine which kinds of statement nodes we can remove, if the corresponding subtree does not contain a NullLiteralExpr node. In the ablation study, we pruned all such subtrees under a particular kind of statement node (and all nodes connected to that subtree via a name layer), for each kind of statement. The resulting F1 scores for the preliminary model appear in fig. 6. Our NaP-AST construction procedure prunes the subtrees and name-connected nodes for the types of statement nodes that appear to the left of the “¡random¿” entry (in red) in fig. 6: that is, those statement types that are less useful to the model than a random statement. For example, a nullability NaP-AST will not contain a for loop (ForStmt in fig. 6) nor nodes for any of the variables manipulated in that loop unless there is a null literal within the loop. However, all expression statements (not within the subtree of another pruned statement) will be preserved, whether or not the ExpressionStmt node for the statement has a null literal in one of its subtrees—removing expression statements damages model performance too much.
ForStmtBlockStmtLocalClassDeclStmtSynchronizedStmtForEachStmtCatchClauseWhileStmtLabeledStmtAssertStmtThrowStmt¡random¿ExplCtorInvocStmtReturnStmtIfStmtContinueStmtSwitchStmtTryStmtExpressionStmt0.80.80.850.850.90.90.950.95Node TypeF1 after dropping
Figure 6. Results of the ablation study in section 3.2.5 by statement type. Subtrees of nodes of types to the left of “¡random¿”, and nodes connected to those subtrees by the name layer, are pruned from the NaP-AST.
5.1.4. Post-processing
Four nullability-specific post-processing steps ensure internal consistency.
Returning Nullable Fields
Our first post-processing step is a simplistic dataflow analysis: when a method returns a field that the model has marked as @Nullable, the method’s return type should be @Nullable, too. This post-processing step is especially important for getters, which the models otherwise rarely predict are nullable.
Consistency between Arguments and Parameters
Our next post-processing step is similar in spirit: at a call-site where an argument is a read of a field that the model predicted is @Nullable, we change the corresponding parameter to @Nullable in the method declaration. We have observed, qualitatively, that these models tend to do a poor job predicting whether parameters are nullable, relative to other kinds of program elements like fields; we conjecture that the reason for this behavior is that nullable fields are often directly assigned null, but many methods with nullable parameters are never directly called with null literal arguments.
Stream Parameters
Since the models use piece-wise boundaries in a continuous, higher dimensional space, their impressionistic understanding of Java syntax occasionally puts annotations where they cause compilation errors. For example, a model might place a @Nullable annotation in a stream parameter, where Java’s syntax does not permit one. We observed this in practice, so we added a post-processing step to remove these @Nullable annotations.
Table 1. GTN hyperparameters
Hyperparameter Value
GT Layers 5
FGTN Layers 2
Epochs 10/batch
Learning Rate 0.01
Channel Agg. Mean
# of Channels 2
Non-local weight -2
k-Hop 9
Inheritance
A method’s inferred parameter or return types in the same inheritance hierarchy may not match. We remove @Nullable annotations in superclasses and interfaces when no such annotation was inferred in the subclass or implementing class, respectively.
6. Training
We split the nodes with @Nullable annotations into 60% training, 20% validation, and 20% test datasets. To create a balanced dataset, we had to combine each set with an equal number of nodes without @Nullable annotations. To do this, we took the annotated nodes in the dataset and subjected them to the 60-20-20 split. We combined the training and validation sets with unannotated nodes of equal number from the same classes, and put all remaining nodes in the test set.
We trained the GTN on a cluster with Intel Xeon E5-2630-v4 CPUs (2.2 GHz), 2 GPUs and 50 GB RAM in the Lochness HPC. Since the adjacency matrices that we use to represent the AST are sparse, they take up a lot of space on the GPU. So, we were only able to train a sample consisting of 1.15% of our dataset at a time. We thus chose batch training, which allows us to train the GTN on the full dataset over a longer period of time. We fixed the input size to 8000 nodes, and only trained on as many classes at a time as fit into that space. The GCN’s training (and all the experiments in section 7) were run on an AWS g4dn.xlarge instance and a Deep Learning PyTorch AMI. We trained NullGCN using PyTorch and the PyTorch Geometric library. The GCN uses two convolutional layers, dropout regularization, negative log-likelihood loss function and the Adam optimizer.
6.0.1. Hyperparameters
After some trial and error, we used a simple 2 layer GCN with batching. The GTN’s trained hyperparameters are in table 1. These hyperparameters were chosen by manual grid search. For this dataset, non-local weight of -2 performs well independent of other variables. Since we trained a large dataset in batches, we chose a low initial learning rate and used weight decay throughout the training process to avoid overfitting.
6.1. Thresholding during Deployment
The models are trained on an equal number of nodes that are nullable and non-null, and all the training data was sourced from classes that contained at least one nullable member. Their predictions reflect this: averaged over many pieces of code, the average nullability prediction is about half nullable elements and half non-null elements. In real code, however, there are many more non-null elements than nullable elements: in fact, real code contains many classes that have no nullable members at all! To account for this, when applying each model, we use a thresholding strategy: we only actually annotate elements as nullable when the model is very confident in that element’s nullability. In our experiments with real projects in section 7.1, we applied a threshold of 90% confidence. In a deployment of one of these models on a previously-unannotated codebase (which is our envisioned deployment scenario), we recommend using a similar threshold.
7. Experiments and Results
We address the research questions from section 1:
  1. RQ1
    What kind of machine-learning model architecture is suitable for inferring pluggable types, and how does that model differ from those in prior type inference works for other languages?
  2. RQ2
    How much data is needed to make that model effective at inferring qualifiers in practice?
We performed two experiments. The first is a large-scale evaluation of each model’s ability to infer ground-truth, human-written annotations on real open-source projects from a previous study of a nullability checker (section 7.1) to answer RQ1. The second experiment (section 7.3) evaluated how much training data the best model from section 7.1 required for good performance, to answer RQ2. All of our code, dataset, and experimental scripts are open-source and available (Kazi Siddiqui, Martin Kellogg, [n. d.]).
7.1. NullAway Benchmark Experiment
We ran each model on the open-source benchmark from the NullAway paper (Banerjee et al., 2019b). This benchmark (from their artifact (Banerjee et al., 2019a)) is a collection of 18 projects from GitHub; each was annotated with nullability annotations when the artifact was published. The artifact also includes an evaluation tool to run NullAway on these projects. Two of these repositories no longer exist; we could build the project and run NullAway on 11 of the remaining 16; two more projects no longer use @Nullable annotations (in the branches pulled by the artifact’s evaluation tool). We consider the remaining 9 projects that we can typecheck using NullAway; however, we note that of these projects, 2 of the 9 do not typecheck with no warnings, even with the human-written annotations. We assume that this is because of either changes in the projects or changes in NullAway itself.
In the remaining projects, 142 out of the 594 classes (~23%) have a @Nullable annotation. This is a typical ratio; as another point of comparison, consider the eureka and spring-framework projects (both of which are outside the dataset but are annotated for NullAway). Eureka  [^5] https://github.com/Netflix/eureka has 21/408 annotated classes (~5%); spring-framework  [^6] https://github.com/spring-projects/spring-framework has 3016/8800 (~34%). These projects are examples of human-annotated projects near the extremes of the distribution.
We checked each benchmark project with NullAway in three configurations: 1) with the human-written annotations that came with the projects, 2) after removing the human-written annotations from each project, and 3) after using each model to re-annotate the code. We measured two proxies for human annotation effort: each model’s precision/recall of the human-written annotations (which are erased before passing the projects to each model) and the reduction in the warnings from NullAway using our annotations versus the unannotated code. The results appear in table 2.
Table 2. NullAway benchmark results. “War.” means “Warnings.” “Red. %” means “reduction percentage”: that is, the reduction in warnings achieved by each model vs. the unannotated code.
Unannotated NullGCN TDG GCN NullGTN GPT-4 CodeQwen
Project War. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec.
ColdSnap 256 2 99% .06 .2 256 0% .72 .28 2 99% .15 .8 0 100% .65 .28 2 99% 1 .26
QRContact 0 0 - 0 0 0 0% .66 1 0 - 0 0 0 - 1 1 0 - 0 0
jib 101 122 -20% .62 .24 101 0% .9 .33 101 0% .55 .68 166 -64% .82 .33 101 0% .9 .28
meal-planner 1 11 -1000% .07 .33 1 0% 1 .66 1 0% .2 .33 9 -800% .25 .66 1 0 1 .66
uLeak 12 6 50% .14 .4 12 0% 0 0 8 33% .41 1 6 50% 1 .2 12 0% 0 0
AutoDispose 2 30 -1400% 0 0 1 50% 0 0 2 0% .01 .11 30 -1400% 0 0 24 -1100% 0 0
picasso 0 0 - .61 .39 0 0% .28 .8 0 - .68 .73 3 - .66 .02 0 - 0 0
skaffold-tools-for-java 1 1 0% 1 .71 1 0 0 0 1 0% .62 .71 2 -100% 0 0 1 0% 0 0
butterknife 4 4 0% 0 0 4 0% 0 0 0 100% 1 1 4 0% 0 0 4 0% 0 0
Totals 377 176 53% .31 .27 376 0% .57 .36 115 69% .39 .69 220 41% .76 .25 145 61% .91 .21
As an example of the the task that the models are being asked to do, consider the following line:
private static LeakInstanceTracker trackerSingleton = null;
A model should infer that this field is @Nullable, and such an annotation is present in the ground-truth human-annotated code; without an @Nullable annotation, NullAway complains that trackerSingleton is supposed to be non-null and yet it is assigned to null. Adding the annotation automatically would reduce the human effort to adapt this codebase to using NullAway, which is why we use warning reduction as a proxy for human effort.
7.2. RQ1
Our experiments compared the following approaches:
  • Graph Convolutional Networks (GCN) (Zhang et al., 2019a) is the common baseline model for node prediction tasks. We trained GCN models on the NaP-AST representation derived in section 5 (“NullGCN”).
  • The TDG encoding from prior work (Peng et al., 2022b) (“TDG GCN”).
  • A Graph Transformer Network (GTN) (Yun et al., 2019) is a model that is mathematically guaranteed to perform at least as well as a GCN, but extends its capabilities for graphs with multiple edge types. Our GTN model (“NullGTN”) is also trained on our NaP-AST encoding; the TDG encoding only has one edge type, so it gains no benefit from the GTN over the GCN.
  • Large Language Models (LLMs) like CodeQwen 1.5-7B-Chat and GPT-4 are a good baseline for coding tasks.
NullGCN
On the NullAway benchmark, NullGCN yields 176 warnings vs 377 unannotated (53% reduction), with precision 0.31 / recall 0.27. While totals improve, the model regresses on some small projects (e.g., meal-planner 1 to 11, -1000%; AutoDispose 2 to 30, -1400%), indicating that misplaced annotations can introduce new warnings even as others are removed. While this is better than no inference at all, this result suggests to us that the hypergraph structure of the GTN (which does much better) is important to modeling where humans write annotations: the same graph is not as useful if the edges are not known to be of different types.
TDG GCN
TDG GCN achieves precision 0.57 / recall 0.39 but yields  0% warning reduction (376 vs 377). The inferences placed by this model were not accurate enough to reduce warnings on the projects. We hypothesize that the TDG chains are too disconnected to place annotations in a consistent manner. We speculate the TDG chains are better suited for use as an additional layer in the NaP-AST rather than an independent encoding. However, because the TDG encoding only has one edge type, it cannot benefit from the GTN model, and it is dramatically outperformed by the NaP-AST plus GTN combination. This shows that the dataflow hints in our name layer are critical for inferring pluggable types.
NullGTN
We built our GTN model (NullGTN) by modifying a fork of the original code (Yun, 2023). It takes an average of about 16 hours to train the model. The largest input source code we evaluated (6,470 non-comment, non-blank lines) took about 5 minutes for inference. NullGTN achieved an F1 score of .95 on the test set during training. On the NullAway Benchmark, NullGTN discovered 69% of the human-written type qualifiers—most of them—with a precision of 39%. However, this only eliminated 69% of the NullAway warnings that a human would need to resolve to adapt these benchmarks to use NullAway—a large amount, but far from all of the warnings. One source of NullGTN’s inability to reduce warnings further is imprecision, which leads to extra annotations, which leads to extra warnings. Another cause of low warning reduction for NullGTN is interactions with libraries. NullGTN has no model of what libraries do, and unlike an LLM it cannot guess based on e.g., the names of library methods. The jib project (where NullGTN fails to remove many warnings despite good recall) uses libraries extensively. It is a modular system with many APIs and plugins, which the model was unaware of. Overall, these results are encouraging: NullGTN makes a good start on bringing these benchmarks into compliance with NullAway, and finds a majority (69%) of the annotations that the developers themselves keep in their codebases.
class LeakTracker {
private WeakReference<Object> gcIndicatorWeakReference;
LeakTracker(Class<?> eventClass, WeakReferenceGenerator g, String name, LeakType leakType) {
this.eventClass = eventClass;
this.weakReferenceGenerator = g;
this.name = name;
this.leakType = leakType;
Figure 7. Simplified example of code that CodeQwen fails to annotate correctly from uLeak’s LeakTracker.java.
GPT-4
To evaluate GPT-4’s suitability for pluggable type inference, we tested its performance on the same dataset used for our NaP-AST-based models. GPT-4 demonstrated moderate precision (76%) and recall (25%) compared to the graph-based models, indicating that while it can correctly annotate some types, it struggles to identify the majority of annotation sites. Despite slightly higher recall than CodeQwen, GPT-4’s more aggressive edits occasionally introduce warnings (e.g., picasso 0 to 3), which explains its lower net reduction.
GPT-4’s strength lies in its general coding abilities, but its lack of domain-specific understanding of pluggable type systems hinders its performance. Unlike NullGTN, GPT-4 does not incorporate graph-based dataflow hints, which are crucial for accurate type inference.
CodeQwen 1.5-7B-Chat
On the NullAway benchmark (Table 2), CodeQwen achieved 91% precision and 21% recall, reducing warnings by 61% overall. However, its edits can increase warnings on some projects (e.g., AutoDispose: 2 to 24 warnings), consistent with occasional unstable edits even at temperature 0.
The precision value is very high because it places relatively few annotations compared to the other models. This is because of CodeQwen’s greater conservatism in making changes. For example, in the code example in fig. 7, the gcIndicatorWeakReference field should be nullable because it is not initialized by the constructor. But since LLMs are trained to reason based on present tokens instead of absent ones, CodeQwen misses it.
Answer to RQ1
Across the 9-project benchmark (Table 2), NullGTN attains the highest recall (0.69) and warning reduction (69%) at moderate precision (0.39). CodeQwen has the highest precision (0.91) but lower recall (0.21) and reduces warnings by 61% overall. The other models and encodings all perform strictly worse than one of these two models. Graph models—especially NaP-AST + GTN—deliver the best overall downstream performance (highest recall and largest warning reduction). LLMs achieve very high precision (e.g., CodeQwen 91%) but lower recall. This suggests that a graph encoding is the right choice for type inference, which is in line with prior work on non-pluggable type inference (Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021).
1020304050607080901000200200400400600600Training percentageWarnings
Figure 8. Warnings by percentage of data trained.
7.3. RQ2
We trained the GTN model with random samples of the dataset at 10% increments (10%, 20%, …) and ran the evaluation to see how much data is required for good results. We wanted to present the evaluation of the downstream task instead of presenting standard ML metrics (like F1 score) on their own, which can be potentially misleading: even a classifier that seems to perform well might not be useful for the downstream task.
Figure 8 shows the result: a classic diagram of the model improving with training data until 60%; performance then worsens after 70% with overfitting. Performance is maximized from training around 50% of the dataset: around 16,000 classes. This estimate gives us a recommendation to calibrate the dataset size for similar tasks.
7.3.1. Availability of Data for Other Pluggable Type Systems
We collected annotations using Sourcegraph’s (Slack and Liu, 2013) src tool to investigate whether any other pluggable type systems (for Java) besides nullability have enough publicly-available data to plausibly train them the same way we trained NullGTN. We found that none did. The closest were @GuardedBy (from the Checker Framework’s Lock Checker (Ernst et al., 2016)) with 1,752 annotated classes; and integer @NonNegative and @Positive annotations (from several sources, including the Checker Framework’s Index Checker (Kellogg et al., 2018)) with 1,675 and 1,008 annotated classes, respectively. These occur an order of magnitude less often than @Nullable, showing there is insufficient data for us to repeat our experiments with another pluggable type system until humans annotate more classes: a “chicken-and-egg problem” for type qualifier inference.
7.3.2. Discussion: “Chicken-and-Egg Problem” for Type Qualifier Inference
The present work does not try to solve this “chicken-and-egg problem” of needing data to enable type qualifier inference (as proposed in this paper), while human adoption is needed to produce that same data. Our contribution is to show that simple feature engineering and choice of model is sufficient to do a reasonable job: that is, we have shown the feasibility of our approach, if enough data is available. However, our work does give the community a sense of how much “human adoption” is necessary before training an inference model like NullGTN is practical, which we believe is a useful step towards an otherwise intractable problem. Further, our work shows that once the threshold for ML-based qualifier inference is reached, it can be leveraged to help increase adoption more.
8. Limitations and Threats to Validity
8.1. Threats to the Validity of Our Experiments
One threat to our NullAway experiment (section 7.1) is that the quality of the human annotations that we used as ground truth is somewhat questionable: they do not remove all NullAway warnings on 3 of the projects, which may distort our results. Another threat is the use of proxies for human effort: precision/recall of human-written annotations and reduction in warning count may both be poor proxies for effort. For example, even though NullGTN finds most of the annotations a human would write, those it fails to find might be the most difficult or complex to write. Another threat comes from the choice of benchmarks: we use programs that have been, at some point, annotated by a human to typecheck with NullAway, so that we have ground truth for our recall computation. However, in a deployment of NullGTN, we expect that the program would never have been human-annotated. Programs which have never been annotated might be trickier for inference, since human annotators might also change code to simplify the process of annotation (and those changes might stay in the code, even if the annotations themselves are later removed). The use of NullAway itself is also a threat to generalizability: our approach may not be as effective at inferring annotations that are useful for some other nullness analysis, such as Meta’s NullSafe.
Our approach may not generalize beyond Java or nullability. In theory, our approach could be applied to other languages beyond Java, but evaluating that is out-of-scope. We expect that the NaP-AST construction procedure in section 3.2 and a FastGTN architecture will be applicable to other pluggable typecheckers that prove the absence of other problems in Java, if a dataset of type qualifiers for those checkers were available. However, we cannot test whether this is effective in practice, since as we showed in section 7.3.1 sufficient training data does not exist for typecheckers that target problems other than nullability: there is at least an order of magnitude more code annotated for nullability than for any other typechecker in open-source, and in section 7.3 we showed that NullGTN required about half of all the data we collected before it became effective.
Another threat to generalizability is that the data experiment in section 7.3 may not be accurate for other pluggable type systems. In particular, nullability type systems, compared to other extant pluggable type systems, are both relatively simple (only two type qualifiers, in a relatively simple type hierarchy) and relatively annotation-intensive (because null is pervasive in most Java programs). We believe that both of these features probably bias nullability inference to require less data, so our estimate may be overly conservative—type systems with large type hierarchies, dependent types, or other complex features may require much more data than our estimates suggest.
8.2. Generality
Our approach works for pluggable type systems with a finite set of types, assuming sufficient training data. The restriction to a finite set of types is an important limitation: we must have fixed the set of type qualifiers to place before training the model, because we’re using a classifier. Many pluggable type systems, such as units-of-measure (Xiang et al., 2020) or integer signedness (e.g., @NonNegative and @Positive in section 7.3.1), fit this model. However, systems with parameterized qualifiers, like @GuardedBy (section 7.3.1), require a different approach—possibly combining our method (to predict annotations) with a generative model (to generate parameters).
In practice, human-annotated code is an order of magnitude rarer for other checkers than the amount needed to train NullGTN (shown in section 7.3.1). This prevents us from directly evaluating the generality of our approach. We’re investigating approaches to synthesize training data for other pluggable type systems, but that is another, separate problem.
8.3. Other Limitations of Our Approach
All machine-learning-based techniques share one serious limitation: there are no guarantees that the annotations that it infers are correct. In our experiments, NullGTN demonstrated close alignment with human-written ground truth annotations—with a recall of 69%. Though there is also no guarantee that human-written annotations are correct, we believe that NullGTN’s ability to hew so closely to developers’ annotations suggests that this problem will be minor in a practical deployment: whether a human or NullGTN writes an incorrect annotation, a sound typechecker will reject it.
9. Related Work
The most closely-related work to ours is a recent approach for “error-driven” inference of nullability annotations for pluggable typecheckers (Karimipour et al., 2023). The key idea of this approach is to treat the typechecker as a “black box” and use its error messages to drive inference: for example, if the checker warns that a return expression is nullable but that a return type is declared as nonnull, this approach could infer that an @Nullable annotation should logically be placed on the return type. In experiments using NullAway, this approach was shown to reduce the number of warnings emitted by the checker by 69.5% (vs. our 60%); they did not measure precision/recall of ground-truth annotations. However, our techniques are different: they relies on the checker’s output, while we rely on where humans have written nullability annotations historically (i.e., the training data for our model). In methodology, the two approaches are orthogonal; we hope to profitably combine them in future work, since the differing approaches mean that they probably infer different subsets of nullability annotations. Another closely related approach bootstraps global inference from the local inference that is built into most pluggable typechecking frameworks (Kellogg et al., 2023). This approach, though not specific to nullness, also uses a fundamentally different approach and might be profitably combined with ours in future work: it works deductively, unlike our approach.
Some prior works (Xiang et al., 2020; Li, 2017) use constraint-based type inference for pluggable types. These approaches do not support nullness (they are specific to a units-of-measurement type system), and cannot be readily applied to other type systems: each is unique. Our methodology, by contrast, could be applied to any type system, if training data were available: a model developer would need to repeat our ablation experiments to drop appropriate nodes from the NaP-AST, but otherwise could reuse our infrastructure. Prior approaches for explaining type inference errors (Lerner et al., 2007; Pavlinovic et al., 2014; Loncaric et al., 2016; Zhang et al., 2017) and for gradual type migration (Phipps-Costin et al., 2021) suffer the same problems as other constraint-based approaches (i.e., do not apply easily to new type systems). There is also prior work that involves the user in type inference (Vakilian et al., 2015), which is complementary to ours. The Daikon invariant detector (Ernst et al., 2007) can insert @Nullable annotations at program locations that can be null at run time (The Daikon developers, 2023). Like any dynamic approach, it is inherently limited by the quality of program’s test suite.
There is a lot of work on inference of static types for dynamically-typed languages such as Python or JavaScript using machine learning (Xu et al., 2016; Pradel et al., 2020; Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021). These approaches are not directly applicable to our problem, since they do not reason about nullability, but about the base type system, and their models and encodings are specialized to that problem.
There is a large literature on type inference and type reconstruction algorithms based on constraint solving (Pierce, 2002, Chapter 22). These techniques build and then solve a system of constraints; the specific constraints are type-system-dependent, which makes this approach unsuitable for pluggable types: determining the constraints to solve would require translating the typechecker’s logic into the logic of the constraint solver, which effectively would require re-implementing the typechecker. Recent work (Madsen and van de Pol, 2021) has proposed an extension to a classic constraint-based type inference approach (Damas and Milner, 1982) that enables reasoning about nullability. However, extending this to the mainstream object-oriented languages (and their associated nullability checkers) is an open-problem.
10. Data Availability
The data and code from our experiments are open-source and available (Kazi Siddiqui, Martin Kellogg, [n. d.]).